### v1

# WiLLM: An Open Wireless LLM Communication System

Abstract.
The rapid evolution of Large Language Models (LLMs) threatens to overwhelm the existing wireless infrastructure, which calls for architectural innovations and frameworks to support the surging mobile LLM services. This paper proposes WiLLM, the first open-source wireless communication system specifically designed for LLM services, consisting of three progressive innovations. Firstly, we establish a novel paradigm by deploying LLMs in core networks (CNs) with abundant GPUs to enable distributed LLM inference services, strategically positioning LLM inference at the convergence point of the backbone bandwidth and the edges of the cellular network. Secondly, we propose an innovative extension to the conventional network slicing architecture, redefining its design with a “Tree-Branch-Fruit” framework. Its specialized adaptation for LLM services enables telecom operators to monetize these services through slice subscriptions while retaining infrastructure ownership. Finally, to operationalize this architectural vision, we present WiLLM, which systematically addresses critical limitations in existing solutions through novel key capabilities previously unattainable: (1) enhanced slice orchestration, extending existing primary slicing to implement a dual-layer slicing architecture and achieving coordinated multi-UE-multi-slice scheduling mechanisms that provide finer-grained resource allocation for LLM services; (2) universal UE compatibility through application-layer tunneling mechanisms, enabling legacy devices without native slicing capabilities to utilize LLM slice services without hardware upgrades; (3) dual-mode scheduling and cross-layer APIs, supporting flexible deployment from CNs to servers. Built on OpenAirInterface, WiLLM extends this established framework with innovations specific to LLM services, lowering the barrier for adoption by researchers without wireless expertise. Furthermore, we release the first LLM wireless communication dataset with 1,649,996 records and synchronized 58-dimensional metrics, along with two relevant benchmarks. A small but quintessential case study of smart glasses demonstrates practical viability in resource-constrained devices and human-machine interaction scenarios. WiLLM aims to build an open platform and ecosystem for cross-layer optimization and AI-telecom convergence. The code, datasets, hardwares and videos are available at https://openwillm.github.io.
Large Language Model, Wireless Communication System
copyright: none
Figure 1. Contributions and advantages of the proposed WiLLM.
1. Introduction
The convergence of artificial intelligence and wireless communications represents one of the most significant technological intersections of the current decade. As computational capabilities and algorithmic sophistication continue to advance, LLMs have emerged as transformative tools with far-reaching implications across diverse application domains. The integration of these models into wireless communication systems has catalyzed a paradigm shift, enabling advanced intelligent services ranging from real-time multilingual translation to context-aware augmented reality applications (Bariah et al., 2024). As this integration accelerates, telecommunication infrastructure faces unprecedented challenges in supporting the computational demands of these resource-intensive models. Current deployments predominantly rely on either edge computing paradigms or centralized cloud processing approaches, each presenting significant limitations. Edge deployment minimizes transmission latency but suffers from constrained computational capacity and thermal limitations in user equipment (UE), while cloud-based solutions introduce unpredictable internet latency that severely impacts quality-of-experience (QoE) for interactive LLM services (Zheng et al., 2024). Moreover, the exponential growth in concurrent LLM service requests threatens to overwhelm both communication channels and computational resources, creating critical bottlenecks in service provisioning.
Recent implementations across several institutions in China, including deployments of Deepseek at multiple universities (Engineer, 2025), demonstrate the feasibility of local LLM deployments at the network edge (Boateng et al., 2024). While these edge deployments represent a promising step toward LLM integration within telecommunications infrastructure, they still encounter resource limitations and scalability constraints inherent to edge computing. These implementations nonetheless validate the technical viability of localized LLM services and establish a foundation for our proposed migration of LLM computational resources to the core network, where greater processing capacity and more efficient resource allocation can be leveraged. This transition from edge to core network deployment represents a natural evolution in the LLM service architecture, addressing the scalability limitations of edge deployments while maintaining the latency advantages over cloud-based solutions (Chaoub and Elkotob, 2025).
The current network architecture exhibits three fundamental drawbacks hindering effective LLM service integration: (1) the absence of dedicated open testbeds for LLM-specific communication scenarios, (2) the inherent conflict between centralized computation paradigms and distributed network architectures, and (3) the lack of viable business models for telecom operators in LLM service provisioning. Existing network slicing architectures provide potential foundations for addressing these challenges through resource isolation and service customization. However, conventional slicing methodologies primarily focus on traditional communication services and lack the specialized mechanisms required for computationally intensive LLM workloads (Dehkordi et al., 2023). Furthermore, the absence of standardized APIs and resource management frameworks inhibits the development of cross-layer optimization strategies essential for efficient LLM service delivery.
Another critical limitation is the high usability barrier of current systems. Existing platforms require specialized wireless expertise and native slicing support on user equipment, restricting access primarily to communication specialists (Zheng et al., 2024). This complexity explains why testbeds are often used only by their developer communities, significantly limiting cross-disciplinary research and broader adoption for LLM service experimentation. Moreover, the field lacks comprehensive empirical datasets capturing cross-layer performance metrics, standardized benchmarking methodologies for systematic comparison, and representative case studies demonstrating viability in resource-constrained environments. This empirical deficit perpetuates the disconnect between algorithmic innovation and operational deployment, hampering the development of generalizable principles for wireless LLM communication systems.
Motivated by the above system and networking challenges, we present WiLLM, the world’s first open-source wireless communication system specifically designed for mobile LLM services. Building upon the solid foundation provided by OpenAirInterface’s modular architecture, WiLLM strategically extends and enhances this established framework with specialized components for LLM service integration, creating a system that maintains compatibility with existing infrastructure while introducing transformative capabilities for AI-driven communications research.
The work unifies the following progressive contributions:
1.1. Paradigm
Novel Network Paradigm for LLM Services (Section 3). WiLLM establishes a novel paradigm by deploying LLMs in core networks (CNs) with GPU infrastructure, enabling distributed LLM inference services strategically positioned at the convergence point between high-bandwidth backbone connections and cellular network edges. This architectural configuration optimizes computational-communication tradeoffs while facilitating operator monetization through subscription-based network slice allocation.
1.2. Architecture
“Tree-Branch-Fruit” Slicing Architecture (Section 3.3). We propose a multi-layered slicing architecture wherein the “tree” represents foundational network infrastructure, the “branches” correspond to conventional communication service slices, and the “fruits” denote specialized LLM service slices. This architecture establishes a “slice-as-LLM-service” paradigm while maintaining compatibility with existing slicing mechanisms, achieving coordinated scheduling across multiple users and slices through cross-layer resource optimization.
1.3. System
The Wireless Communication System Specialized for LLM Service (Section 4). Based on OpenAirInterface, WiLLM implements five pivotal innovations: dynamic slice compatibility, universal UE compatibility, dual-mode resource scheduling, cross-layer API framework, and flexible deployment support. This integrated solution addresses critical limitations in existing wireless infrastructure while preserving backward compatibility.
1.4. Dataset
The World’s First LLM Wireless Communication Dataset and Benchmarks (Section 5). We introduce the first comprehensive LLM communication dataset with 1,649,996 records, each containing 58 communication metrics synchronized across UE-gNB-CN interfaces. This enables holistic cross-layer analysis of LLM communication patterns. In addition, we present two metrics, LAREI and LSEQ, as benchmarks that offer a standardized methodology for wireless LLM service evaluation.
1.5. Case Study
Smart Glasses Case Study (Section 6). The practical utility of WiLLM is validated through a smart glasses application that implements gesture-triggered queries and scene interpretation. Based on the principle of user-perceived continuity in human-computer interaction, we establish a response latency target of approximately 2 seconds, achieving optimal slicing through offline statistical analysis and online learning methodologies.
2. Related Work and Comparison
2.1. Wireless Communication for LLM
The convergence of LLMs with wireless systems has precipitated prolific theoretical advancements characterized by a critical implementation deficit. Despite many approaches being proposed, the field confronts a fundamental disjunction between algorithmic innovation and operational realization.
Architectural paradigms manifest in three configurations: 1. Cloud-edge frameworks (Chen et al., 2024b) (Wu et al., 2020) (Dong et al., 2024) establish a hierarchical computational distribution by allocating processing tasks between the cloud and edge devices. This approach combines the high computational power of the cloud with the low latency characteristics of the edge. For example, NetGPT (Chen et al., 2024b) relies on cloud-edge collaboration in personalized generative services. However, the dynamic nature of wireless channels limits its adaptability. Reference (Wu et al., 2020) explores the potential of cloud-edge computing, while (Dong et al., 2024) highlights the challenges of implementing edge AI generated from cloud-based LLMs. 2. Edge computing deployment (Cai et al., 2024) (Nelson and Toloui, 2023) (Tang et al., 2023) runs LLMs directly on edge servers or devices, enhancing performance through quantization techniques. For instance, Edge-LLM (Cai et al., 2024) optimizes edge adaptability using unified compression and adaptive layer voting, while (Nelson and Toloui, 2023) demonstrates hardware-supported edge deployment cases. Nevertheless, the limited capacity of edge devices remains a bottleneck, as further validated in (Tang et al., 2023). 3. Distributed architectures (Xue et al., 2024) (Mitra et al., 2024) decompose LLM computational tasks across multiple nodes to fully utilize distributed resources. For example, WDMoE (Xue et al., 2024) proposes a wireless distributed framework based on mixture-of-experts, though its validation in congested scenarios remains insufficient.
Resource optimization methodologies span cross-modal approaches (JPPO (You et al., 2024)) integrating prompt compression with power optimization; adaptive partitioning strategies (Chen et al., 2024a) reconfiguring computational boundaries under network fluctuations; and collaborative frameworks (He et al., 2024) optimizing offloading decisions across distributed resources. Service integration architectures encompass specialized slicing methodologies (LLM-Slice (Liu et al., 2024)), intelligent orchestration approaches (Dandoush et al., 2024), and long-context integration strategies (Xu et al., 2025; Zhang et al., 2025).
Bottlenecks and our solutions: The field faces three bottlenecks: theory-centric validation lacking real-world relevance, insufficient cross-layer integration, and unproven business viability. WiLLM addresses these through an empirical platform bridging theory and implementation. Unlike conventional edge server strategies that incur coordination inefficiencies, we deploy LLMs on GPU-equipped infrastructure at the backhaul-RAN nexus, enhancing computational capacity, resource optimization, and infrastructure integration.
2.2. Related Systems and Comparisons
In this section, we evaluate prominent open-source systems for their capabilities across a range of technical features essential for modern wireless networks, including OAI, srsRAN, Open5GS, FlexRIC, and Janus. While these systems contribute significantly to wireless network research and development, they exhibit limitations in fully supporting the comprehensive demands of next-generation networks. OAI provides a complete 4G/5G protocol stack implementation, spanning from the physical layer to the network layer. Its modular design supports academic research and prototyping (Nikaein et al., 2014). srsRAN is a lightweight and highly customizable 4G/5G software suite (Gomez-Miguelez et al., 2016), offering flexibility in scheduling and QoS management. Open5GS implements comprehensive 5G core network functions, including AMF, SMF, and UPF components (Barbosa et al., 2024), establishing a solid foundation for network integration. FlexRIC is a programmable RAN SDK focused on 5G multi-tenancy and low-latency optimization (Schmidt et al., 2021). Janus is a programmable RAN platform developed by Microsoft, based on the O-RAN architecture, supporting dynamic service models (Foukas et al., 2023). While these systems offer essential assistance in certain aspects of wireless networking, they exhibit inadequate support for general technical features like UE compatibility, slice flexibility, and specialized support for LLM communications, etc. Detailed objective metric comparisons are presented in Table 2 and Table 1 in the Appendix.
2.2.1. Research Gaps and Positioning of WiLLM
While existing open-source communication platforms provide valuable research infrastructure, comprehensive analysis reveals significant operability impediments that restrict their adoption beyond specialized research communities. As quantified in Table  1 and 2, current systems demonstrate substantial limitations in critical usability dimensions. The requirement for native slicing support at the UE level presents a particularly problematic constraint, effectively excluding standard commercial devices without specialized protocol modifications (Saboorian and Xiang, 2017; Andrade and Wickboldt, 2025). This limitation fundamentally contradicts the purpose of experimental testbeds, which should facilitate algorithm validation across diverse implementation scenarios rather than constraining experimental configurations to narrow protocol-specific deployments. Furthermore, these resulted in an undesirable ecosystem fragmentation where testbeds remain utilized primarily by their own development communities, significantly limiting cross-disciplinary innovation.
WiLLM aims to bridge the technological gap between wireless communication and LLM systems by extending the OAI architecture to build the first open-source wireless communication system designed specifically for LLM services.
3. Paradigm
Figure 2. Left: Conventional cloud-based LLM service deployment architecture with internet connectivity. Right: Proposed core network LLM service deployment architecture with GPU resources at network convergence points.
LLMs as transformative technologies in wireless communications call for fundamental network architecture reconceptualization. Despite edge computing’s partial success in LLM deployment, its device-centric approach inherently limits scalability and optimization. We propose a novel core network paradigm that positions LLM resources at the nexus of high-capacity backhaul and wireless access networks, enabling comprehensive computational-communication resource integration that transcends traditional edge computing constraints.
3.1. Deployment Paradigm for LLM Services
Current LLM service deployment architectures exhibit fundamental architectural limitations that constrain their performance envelopes. As illustrated in Figure 2, traditional cloud-based configurations necessitate traversal of multiple network boundaries: from UE through gNB and CN to remote LLM servers via public internet infrastructure. This architectural paradigm introduces non-deterministic network latency due to Internet Gateway traversal, unpredictable congestion during peak utilization periods, and critically, a decoupling of computational and communication resource management planes that prevents cross-layer optimization.
[…]
The technical implementation comprises three principal architectural components:
  1. (1)
    Deterministic Service Path Integration: As shown in Figure 2, our implementation establishes a direct service flow from UE through gNB to the CN-GPU integrated infrastructure. This configuration eliminates the indeterministic routing and potential congestion points inherent in public internet traversal, enabling guaranteed service-level agreement enforcement through contiguous administrative control of the entire transmission path.
  2. (2)
    Hierarchical Slice-Based Resource Orchestration: The implementation leverages the slice management mechanisms detailed in Section 3.3 to establish fine-grained coordination between communication and computational resources. Rather than treating these as separate resource domains (as in conventional implementations), our architecture implements a unified resource allocation framework where slice policies govern both radio resource allocation at the gNB and computational resource distribution at the CN-GPU nexus.
  3. (3)
    End-to-End Service Flow Management: The architecture implements a comprehensive flow control mechanism that extends from UE request initiation through transmission scheduling, computational resource allocation, and response delivery. This coordinated approach enables advanced optimization techniques such as predictive resource allocation, where the system anticipates computational requirements based on request characteristics and proactively configures both network and computational resources accordingly.
3.2. Transcending Edge Computing Limitations
The core network deployment paradigm for LLM services introduces fundamental architectural, operational, and infrastructural enhancements that transcend edge computing’s intrinsic limitations. The hierarchical “Tree-Branch-Fruit” slicing framework (Section 3.3) enables differentiated, SLA-governed LLM orchestration, surpassing edge computing’s localized resource allocation. The cross-layer optimization framework (Sections 4.2) synergistically coordinates resources across network strata, realizing dynamic adaptation unattainable through edge computing’s device-centricity. Strategic LLM resource placement at the backhaul-RAN confluence forges a computational-communication symbiosis, transcending edge deployments’ latency-processing tradeoffs. Empirical insights from the WiLLM dataset (Sections 5) corroborate the core network paradigm’s transformative potential, illuminating performance benefits and optimization opportunities eclipsing edge computing. This paradigm establishes the foundation for the “Tree-Branch-Fruit” slicing architecture, facilitating a technical environment where both communication resources and computational capabilities can be allocated through unified slice policies. The implementation enables an evolutionary deployment path that preserves investment in existing infrastructure while advancing sophisticated LLM service delivery models.
3.3. “Tree-Branch-Fruit” Slicing Architecture
Figure 3. Tree-Branch-Fruit network slicing architecture with hierarchical service organization.
Figure 4. Slice function architecture of WiLLM with integrated UE, gNB, Core Network, and Edge Server Components (Liu et al., 2024).
Building upon our core network deployment paradigm, we propose a novel “Tree-Branch-Fruit” slicing architecture that transforms conventional network slicing approaches into a comprehensive service provisioning platform. This multi-tier architecture, illustrated in Figure. 3, establishes a hierarchical framework that maintains backward compatibility with existing slicing mechanisms while introducing specialized LLM service slices. In this architecture:
  • Tree (gNB): Represents the foundational radio access network infrastructure, serving as the common entry point for all network traffic and the root of the slicing hierarchy.
  • Branch Slices: Correspond to traditional communication service slices defined in 5G standards, including enhanced Mobile Broadband (eMBB), Ultra-Reliable Low-Latency Communication (URLLC), and massive Machine Type Communication (mMTC). These branch slices are not the focus and implementations of this work. These branches maintain backward compatibility with existing network functions.
  • Fruit Slices: Denote specialized LLM service slices that extend from branch slices, each configured for specific application domains and computational requirements. Figure. 3 presents illustrative examples of potential fruit slice configurations, including various LLM service use cases with different model sizes tailored to specific application requirements.
The examples illustrated in Figure. 4 demonstrate how different application domains might be served by appropriately sized language models, with computationally intensive applications connected to branches supporting higher bandwidth and more powerful models, while lightweight IoT applications utilize smaller, more efficient models. These examples are not prescriptive implementations but rather illustrative use cases demonstrating the flexibility of the proposed architecture. This architecture establishes a “slice-as-a-LLM-service” paradigm that facilitates efficient multi-user and multi-slice coordinated scheduling through cross-layer resource optimization. The hierarchical structure enables:
  • Differentiated Service Provisioning: Telecom operators can provide LLM services with varying performance guarantees and computational capacities tailored to specific application requirements and price points.
  • Resource Isolation: The architecture enforces strict isolation between service slices, preventing performance degradation in one service from affecting others and enabling guaranteed quality-of-service levels.
  • Domain-Specific Optimization: Each fruit slice can be independently optimized for its specific application domain, with tailored model selection, inference parameters, and resource allocation strategies.
  • Modular Service Evolution: New LLM services can be deployed as additional fruit slices without disrupting existing services, facilitating incremental innovation. Our architecture achieves forward compatibility via application-layer tunneling that transcends conventional slicing constraints. This method encapsulates LLM traffic within standard application protocols, enabling infrastructure compatibility while maintaining architectural autonomy. The approach provides telecommunications operators with implementation flexibility independent of underlying radio network slice capabilities. This strategic decoupling of service implementation from network infrastructure allows fruit slices to evolve asynchronously from hardware upgrade cycles or standardization timelines. Such architectural independence ensures operational resilience across heterogeneous networks, enabling deployment of novel LLM capabilities unfettered by the limitations of conventional slicing frameworks.
The monetary value icons associated with different fruit slices in Figure. 3 illustrate the differentiated pricing potential of this architecture, with higher-resource slices commanding premium pricing while lightweight services remain cost-effective for mass deployment. This economic framework creates sustainable incentives for network operators to invest in LLM-capable infrastructure.
The “Tree-Branch-Fruit” architecture represents an advancement over conventional slicing approaches by introducing application-specific service customization while maintaining backward compatibility with existing network functions. This novel paradigm establishes the foundation for our comprehensive WiLLM implementation, which will be detailed in subsequent sections.
4. The Proposed System
Figure 5. Comprehensive system architecture of WiLLM with UE, gNB, CN, and Edge Server Components.
This section presents the comprehensive design and implementation of WiLLM, our novel wireless communication system specifically optimized for LLM service delivery. Building upon the “Tree-Branch-Fruit” slicing architecture described in Section 3, we have constructed a complete system implementation that extends the OpenAirInterface framework with specialized components for LLM service provisioning. Figure. 5 illustrates the system architecture, encompassing UE, gNB, core network, and edge server components with their associated interconnections and functional modules.
4.1. System Architecture Overview
WiLLM adopts an innovative vertical integration architecture design, achieving end-to-end coordination from user applications to core networks through a hierarchical management framework. This architecture, centered on layered slice management, organically integrates four key subsystems: the CN subsystem built on the Open5GS framework, with enhanced edge computing capabilities through a dedicated Edge Server Bridge; the gNB subsystem that implements enhanced “Tree-Branch-Fruit” slice management mechanisms based on OpenAirInterface; the UE subsystem that integrates configuration management, slice control, and performance measurement function modules; and the edge server subsystem that carries various LLM deployments and executes corresponding slice policies. These subsystems are not simply stacked together but are tightly coupled through a hierarchical slice management framework, enabling precise scheduling and collaborative optimization of cross-layer resources, thereby ensuring the efficiency of the LLM communication system.
4.2. Implementation Framework
WiLLM extends OpenAirInterface with strategic innovations that balance implementation efficiency with transformative functionality. This approach leverages established communication foundations while introducing specialized components that fundamentally transform system capabilities:
4.2.1. Dynamic Slice Compatibility
As illustrated in Figure. 5, our implementation extends conventional network slicing capabilities with LLM-specific service slices while maintaining backward compatibility. The Slice Manager of gNB encompasses both Branch Slice-UE Mapping and Fruit Slice-UE Mapping components, with bidirectional data flows (indicated by the UE Update and UE State Report pathways) enabling dynamic reassignment of resources in response to changing service demands.
The implementation employs a virtualization layer that abstracts underlying slice implementation details, while the Branch Slice Policy Executor and Fruit Slice Policy Executor enforce service-specific policies at the gNB level. As shown in the left section of Figure. 5, these policy executors interact directly with the PRB Manager, creating a decision hierarchy that translates high-level service policies into physical resource allocations.
4.2.2. Universal UE Compatibility
To address the limitations of current UE implementations, which often lack native slicing capabilities, WiLLM incorporates a middleware-based slice access mechanism through the Slice Manager of UE component. The data flow illustrated in Figure. 5 demonstrates how this component interfaces with both the Configuration Manager and Hot Start Module, enabling UEs without native slicing support to access specialized LLM services through application-layer tunneling techniques that encapsulate service traffic within standard data streams.
This compatibility layer effectively bypasses the limitations of current community implementations of primary slicing, which remain incomplete and inconsistently supported across device ecosystems, thereby expanding the addressable user base for LLM services.
4.2.3. Multi-UE Multi-Slice Scheduling Capability
WiLLM implements a comprehensive multi-UE-multi-slice coordinated scheduling mechanism that operates through a two-phase process. The Scheduler module in Figure. 5 forms the nexus of this coordination, receiving inputs from multiple sources including the Buffer Manager (which provides Buffer Status information), HARQ Manager (which signals Retransmission Requirements), and the Slice Allocation of UE component.
The multi-UE-multi-slice coordination operates through a two-phase process:
  • A global allocation phase distributes available physical resource blocks across active slices according to their relative priorities and resource guarantees.
  • An intra-slice scheduling phase allocates resources to individual UEs within each slice based on service-specific metrics.
This hierarchical approach ensures fair and efficient resource utilization while maintaining service isolation, effectively enabling legacy devices to benefit from slice-specific performance guarantees without hardware modifications.
4.2.4. Dual-Mode Resource Scheduling
WiLLM implements complementary “separated” and “embedded” resource scheduling methodologies to accommodate diverse operational requirements and infrastructure constraints:
  • Embedded Scheduling Mode: The embedded approach implements slice-aware resource allocation directly within the scheduler pipeline, as formalized in Algorithm 1. Figure. 5 illustrates how this approach integrates with the Resource Grid and Spectrum Usage Tracking components to maintain awareness of physical layer constraints while implementing policy decisions.
  • Separated Scheduling Mode: The separated approach decouples slice-specific policy decisions from core scheduling logic, employing an external decision engine that solves a global optimization problem and subject to resource constraints and slice isolation requirements. The data flow in Figure. 5 shows how the Resource Update pathway enables this external engine to maintain synchronization with the core scheduling system.
The interaction between the Uplink/Downlink Allocation Request pathways and the PUCCH Allocator modules for both uplink and downlink control channels (as depicted in Figure. 5) demonstrates how the system coordinates the allocation of scarce control channel resources across multiple UEs and slices.
4.2.5. Cross-Layer API Framework
WiLLM introduces a comprehensive API architecture (detailed in Figure. 16) that facilitates vertical integration between application requirements and network-level resource allocation. As shown in Figure. 5, this framework enables multiple feedback pathways:
  1. (1)
    Performance Feedback Loop: The UE State Report, Configuration/Feedback/Buffer, and UE Update connections establish a closed-loop feedback system that continuously monitors service performance and adapts resource allocation accordingly. The bidirectional nature of these connections, clearly depicted in Figure. 5, enables adaptive behavior at multiple timescales.
  2. (2)
    Service-Specific Optimization: The LLM Service ID mechanism enables adaptation of network parameters based on computational characteristics of different LLM implementations. Figure. 5 illustrates how this information flows from the Edge Server Subsystem through the Slice Manager to influence allocation decisions.
  3. (3)
    Resource Utilization Monitoring: The Resource Usage and Slice Allocation connections provide visibility into consumption patterns, establishing a cross-layer information flow that informs both immediate scheduling decisions and longer-term policy refinements.
4.2.6. Deployment Flexibility
The system supports flexible LLM deployment configurations, with the Edge Server Large Language Model Deployment component hosting multiple LLM variants with different parameter sizes as shown in Figure. 5, and the CN-Edge Server Bridge facilitating seamless integration between computational resources and network infrastructure through standardized file and network interfaces.
4.3. The Database Module
A distinctive feature of WiLLM is its integrated data collection module (shown in the upper section of Figure. 5), which continuously refines resource allocation strategies based on operational metrics. This module creates a meta-feedback loop that transcends individual communication sessions, collecting performance data from UE’s Performance, gNB Measurement, and Server Measurement components to inform the evolution of the Customized QoS Scheduler and Model components.
The system incorporates a comprehensive Database component (illustrated in the right section of Figure. 5) that persistently stores operational metrics, enabling longitudinal analysis of system performance. This data repository serves as both an analytical resource for offline optimization and a training corpus for online learning algorithms.
4.4. System Validation Infrastructure
To validate practical performance, we deployed a fully functional testbed (Figure.14 in Appendix) integrating a complete wireless communication chain with LLM service deployment. Following the logical architecture shown in Figure. 5, the core computing infrastructure employs an Intel Core i9-14900KF processor as gNB host, USRP B210 for RF frontend, and NVIDIA RTX 4090 GPU for LLM inference, supporting multi-user, multi-slice concurrent testing scenarios. The system’s operational status can be monitored through a custom GUI (Figure. 15 in Appendix). This configuration achieves complete system integration from radio access to core network, providing an ideal environment for evaluating WiLLM in practical deployments.
5. The WiLLM Dataset
We present the first comprehensive LLM communication dataset with synchronized metrics, enabling reproducible research and systematic evaluation. Appendix F details components, acquisition, synchronization, and error descriptions.
5.1. Dataset Metrics Overview
Our dataset contains 1,649,996 records with 58 metrics spanning the LLM service delivery chain. It covers four scenarios: static UE and gNB (290,653 records), dynamic UE and static gNB (363,906 records), static UE and dynamic gNB (430,369 records), and dynamic UE and gNB (565,068 records). “Dynamic” indicates dynamically allocated network slices, while “static” indicates invariant parameters throughout measurements. The dataset includes 22 metrics at the user equipment layer, 25 at the wireless access network layer, and 18 at the edge server layer.
Figures 17 and 18 confirm stable base station power and noise levels during data collection. We implemented NTP-based distributed synchronization, calibrating all devices through a common server for microsecond-level sampling precision. By recording timestamps at both the UE and server endpoints and using latency compensation algorithms, we maintained synchronization errors within $\pm$1.0 milliseconds.
5.2. Metric Architecture
The dataset provides synchronized metrics across three architectural layers: UE Metrics (Appendix Table 4) capture end-user experience through latency, timing, resolution, modalities, and payload characteristics; gNB Metrics (Appendix Table 6) document wireless conditions via radio parameters (MCS, CQI, BLER), resource allocation, transmission performance, and slice configuration; and CN with Edge Server Metrics (Appendix Table 5) measure computational aspects through inference timing, token processing, model loading, and response generation. The dataset’s value lies in its temporal synchronization, enabling precise correlation between user experience, network conditions, and computational performance. This facilitates advanced research: developing ML models for scheduling, implementing data-driven allocation, creating cross-layer analytics, and validating theoretical models. This corpus establishes a foundation for reproducible research in wireless systems supporting LLM services.
5.3. Preliminary Findings and Conclusions
Figure 6. Component distribution and latency analysis for uplink transmissions across resolution groups. The x-axis labels R1 - R6 represent various resolution groups for image requests sent to LLM, where R1 - R6 correspond to increasing resolutions, with “n” representing the sample size per group, and the lower value of “n” indicating the resolution range for each group. The left panel shows the temporal distribution of communication components (uplink, downlink, inference), the center panel presents statistical trends with standard errors, and the right panel quantifies each component’s percentage contribution to total latency, illustrating how higher resolutions affect time allocation in the transmission pipeline.
Figure 7. Component distribution and latency analysis for downlink transmissions across resolution groups. This figure illustrates the temporal characteristics of downlink transmissions stratified by resolution groupings. Similar to Figure. 6, it provides three complementary views: component distribution, statistical latency metrics, and proportional component analysis. All legend elements in the figure remain consistent with Figure. 6, with the only distinction being the utilization of downlink samples.
This section presents preliminary findings derived from our dataset, with particular focus on communication directionality characteristics and the impact of network slicing on performance metrics.
5.3.1. Uplink vs. Downlink Characteristic Analysis
Finding 1: Our dataset reveals distinct performance profiles for uplink and downlink communication scenarios. In the uplink case (UE image request $\rightarrow$ LLM inference $\rightarrow$ text response), as visualized in Figure. 6, we observe that inference time constitutes a substantial proportion of the total processing latency (74-87% across resolution groups). The relative contribution of uplink transmission time marginally increases (11-25%) with the rise in resolution.
Finding 2: In contrast, the downlink scenarios (text request $\rightarrow$ LLM text response $\rightarrow$ image response), depicted in Figure. 7, demonstrate markedly different characteristics. Downlink transmission time dominates the overall latency budget (81-86%), with inference time contributing a significantly smaller proportion (12-17%).
Conclusions: These directional performance characteristics validate WiLLM’s “Tree-Branch-Fruit” slicing architecture and LLM deployment paradigm in several critical ways:
  1. (1)
    In the uplink scenario, the largest delay comes from inference time, followed by uplink transmission time. This finding supports our strategy of deploying LLMs with GPUs in the CN. This is achieved by reducing the inference time through distributed LLM processing at the CN and minimizing uplink transmission delays by staying close to users, thereby avoiding the need to go through the public Internet.
  2. (2)
    The pronounced asymmetry between uplink and downlink transmission characteristics empirically validates our “Fruit Slice” architectural paradigm, confirming that effective monetization of LLM services requires direction-specific slice configurations that independently optimize bandwidth and computation intensive workloads, a capability that is not achievable through conventional slicing approaches.
  3. (3)
    The pronounced resource requirement asymmetry between uplink and downlink scenarios provides compelling empirical justification for direction-specific scheduling mechanisms. These findings validate the architectural necessity of dual-mode scheduling capabilities, confirming that efficient LLM service delivery requires dynamic, direction-aware resource provisioning rather than single-direction allocation policies. This increased flexibility in provisioning is a key enabler of the slice-based monetization framework.
These insights reveal that WiLLM’s paradigm and architecture, substantiated by real-world datasets, demonstrate critical advantages over conventional systems in operational dimensions. These advancements provide a robust framework for direction-specific resource optimization that traditional wireless systems cannot achieve. We provide further analysis of the reasons for the experimental phenomenon and additional notes in the appendix F.1.
5.3.2. Network Slicing Impact Analysis
Figure 8. Component distribution and latency analysis across slice groups. In the figure, the communication resources for uplink slices in Slice 1 to Slice 3 are gradually increasing. This examines performance impacts of different slice configurations on communication component distribution, highlighting distinct latency profiles for three configurations. The middle panel statistically validates performance differences, while the rightmost panel quantifies how slice configurations shift processing time proportions among uplink, downlink, and inference operations.
Figure 9. This figure presents a temporal analysis of PRB allocation patterns across three implementation scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_rbs”, indicating the current number of resource blocks allocated to the UE. The visualization demonstrates the effectiveness of the slicing architecture in providing differentiated resource allocation while maintaining threshold compliance, with slice-distinguished traffic showing clear separation between service classes. The area below the dashed line is labeled as “basic data” with no data transmission, while the area above corresponds to “core data”.
Figure 10. The figure illustrates a temporal analysis of PRB allocation patterns across three scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom). The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_bytes”, indicating the current transmitted bytes of the UE.
Finding 3: The effectiveness of our “Tree-Branch-Fruit” slicing architecture is demonstrated through comparative analysis of network performance under different slicing configurations. Figure. 8 illustrates how slice configuration affects the temporal distribution of processing components, with notable variations in inference time allocation (43.1% to 59.6%) and uplink transmission time (28.9% to 54.7%) across different slicing arrangements. This demonstrates that slicing indeed affects the latency of LLM communication services and its components, which supports WiLLM’s paradigm.
Finding 4: Examining PRB allocation patterns in Figure. 9 alongside actual byte transmission data in Figure. 10 reveals critical insights into slicing efficiency. Our implementation successfully achieves differentiated PRB allocation conforming to slice policies. Moreover, there exists a discrepancy between theoretical resource allocation and actual transmission patterns. Specifically, byte traffic does not scale proportionally with PRB allocation, suggesting potential resource utilization inefficiencies within the slicing framework. This indicates that due to the complex interplay of elements affecting communication, system execution mechanisms, and varying UE requests, PRB and actual communication bytes are not strictly linearly correlated.
Conclusions: These empirical observations demonstrate the critical value of WiLLM in several dimensions:
  1. (1)
    The significant variability in component allocation across slice configurations validates that WiLLM’s “Tree-Branch-Fruit” architecture enables effective LLM-specific service slicing, supporting our proposed diverse pricing patterns and underscoring the necessity of finer-grained resource allocation mechanisms.
  2. (2)
    The observed allocation-utilization gap reveals that theoretical and simulation-based approaches often fail to capture real-world complexities, confirming WiLLM’s role as an essential empirical platform bridging theory and implementation.
  3. (3)
    The non-linear relationship between resource allocation and actual transmission performance underscores the necessity of WiLLM’s cross-layer API framework and dual-mode scheduling capabilities, which enable adaptive optimization based on observed rather than simulated performance patterns.
The analysis of the causes of the observed phenomenon is presented in the Appendix F.2.
These findings substantiate WiLLM’s fundamental contribution: while traditional approaches rely on simplified resource allocation models, our system provides the empirical foundation and architectural flexibility required for truly efficient LLM service delivery in wireless environments. This empirical grounding enables telecommunications operators to achieve differentiated service quality and optimal resource utilization, a critical requirement for commercially viable LLM wireless services.
5.4. Benchmark
Analyzing datasets reveals intricate dependencies among communication resources, computational capacity, and service quality, requiring specialized metrics. As traditional network metrics fail to capture LLM service delivery’s multidimensional nature, we propose two complementary metrics for evaluating LLM wireless communication systems. The LLM-Aware Resource Efficiency Index (LAREI) quantifies the efficiency relationship between request complexity, computational demands, and resource utilization:
$\text{LAREI}=\frac{\text{RDV}\cdot\log(1+\text{LLM\_Para})}{\text{Resources}\cdot\text{Latency}}\cdot\omega$
Where RDV is request data volume, LLM_Para represents model complexity in billions of parameters, Resources denotes allocated communication resources, Latency indicates end-to-end response time, and $\omega$ is a normalization coefficient.
For slicing effectiveness evaluation, we introduce the LLM Slice Efficiency Quotient (LSEQ):
LSEQ=RDV_slice(1Error Rate)LLM_Para_sliceSlice $\text{LSEQ}=\frac{\text{RDV\_slice}\cdot(1-\text{Error Rate})\cdot\sqrt{\text{LLM\_Para\_slice}}}{\text{Slice Resources}}\cdot\delta$
[…]
These metrics establish a framework for evaluating and optimizing wireless LLM communication systems, balancing computational sophistication with communication efficiency. Detailed analysis is presented in Appendix G.
6. A Case Study of Smart Glasses
6.1. Smart Glasses Integrated in WiLLM
Figure 11. Smart glasses hardware integrates with the WiLLM wireless LLM service, following a typical UE-gNB-CN network architecture. The UE consists of customizable smart glasses and a Mini PC, which also houses the 5G module. A camera positioned at the glabella captures images, enabling collection and querying through a custom human-computer interaction logic.
Figure 12. Gesture-triggered 5G LLM smart glasses data flow. The smart glasses application activates through embedded gesture recognition as the trigger. Upon detecting a five-finger extension followed by a grasping motion, the system captures the current visual feed and transmits it as a request to the LLM.
To validate WiLLM’s effectiveness in real-world applications, we implemented a small but quintessential case study using augmented reality smart glasses as a resource-constrained terminal device requiring LLM integration. Figure. 11 illustrates the experimental setup, showing the hardware components of our custom smart glasses platform interfaced with WiLLM network infrastructure. The glasses feature a 5G module, computing unit, and camera capable of capturing visual data for LLM interpretation. This system enables users to query about scenes before them, allowing the LLM to interpret the visual context through a LLaVA model deployed in the core network host, accessed via gNB.
Key Validation: The smart glasses implementation demonstrates a concrete manifestation of WiLLM’s “Tree-Branch-Fruit” slicing architecture, where the device functions primarily as a visual I/O interface while the computational workload is handled by GPU resources positioned at the network convergence point, as proposed in Section 3.
Figure. 12 shows the interactive paradigm between user and LLM service. The system uses gesture recognition to convert physical movements into contextual prompts for LLM processing, creating a bidirectional channel where gestures initiate queries and responses are displayed through the glasses’ interface.
6.2. Objective from an HCI Perspective
Human-computer interaction (HCI) research shows that interaction quality depends less on minimal latency than on consistent, predictable response patterns (Nabiyouni et al., 2017; Jacobs et al., 1997). Users adapt to predictable systems, but unpredictable delays disrupt interaction continuity. Our goal is to identify configurations that stabilize latency around 2 seconds for the smart glasses, balancing computational efficiency and perceptual continuity within wireless LLM constraints.
Empirical Insight: This case study revealed cross-layer dependencies where UE requests influenced communication resource utilization and allocation at the gNB, which affected LLM processing at the core network. This substantiates the complex relationships between resource allocation and performance documented in Section 5.
6.3. Approaches to Slice Optimization
Figure 13. Online learning progress of the slice selection task for the smart glasses. Gradually converge to the target slice.
We deployed two complementary methodologies to determine optimal slice configuration. Our offline approach utilized WiLLM’s data collection capabilities to analyze six slice configurations, identifying those that consistently maintained the 2-second latency target with minimal variance. Concurrently, we implemented an adaptive online learning framework using the UCB algorithm that continuously monitored performance metrics and dynamically adjusted slice selection. As shown in Figure. 13, this method demonstrated progressive refinement toward optimal configuration with reduced latency fluctuations after convergence. While offline methods provided thorough validation requiring extensive initial data, online learning enabled adaptive optimization with initial exploration costs. Both approaches effectively identified suitable configurations, demonstrating WiLLM’s cross-layer optimization capabilities for resource-constrained applications.
Empirical Insight: The smart glasses case study provides empirical evidence that LLM service quality in wireless environments, particularly in some wearable HCI applications, depends more on maintaining predictable latency through intelligent slicing rather simply maximizing resources. While this single case offers only a directional indicator for resource-constrained applications, it demonstrates that WiLLM’s architectural principles effectively support interactive LLM experiences on computationally limited devices. Our dual optimization methodologies illustrate WiLLM’s availability of the dataset and cross-layer optimization potential, enabling sophisticated resource allocation tailored to specific application requirements. The hardware configurations and software components have been open-sourced to facilitate replication and expansion into other application areas.
7. Conclusion
WiLLM is the first open-source wireless system for LLM services. Our “Tree-Branch-Fruit” slicing architecture, LLMs deployment in CNs, and cross-layer APIs address limitations in existing wireless infrastructure for LLM delivery. The implementation extends OpenAirInterface with components enabling dynamic resource allocation, universal device compatibility, and multi-slice scheduling while preserving backward compatibility. A dataset of 1.6+ million synchronized metrics provides empirical foundation for research, while our smart glasses case demonstrates viability in resource-constrained environments. WiLLM connects telecom evolution with AI, enabling wireless LLM service chain optimization. Future research efforts will focus on addressing the current system’s deficiencies by prioritizing improvements in system stability and fault recovery, ensuring strict compliance with protocols, and implementing security measures for cross-layer APIs.
Appendix
Appendix A Appendix Overview
This appendix provides supplementary information, analysis, and technical details to support the main body of the paper. The contents are organized as follows: Appendix B presents a general comparison of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks. This comparative analysis establishes WiLLM’s distinctive contributions within the telecommunications-AI convergence landscape.
Appendix C extends the comparative analysis through a granular examination of specific technical capabilities across contemporary wireless systems. This detailed evaluation substantiates WiLLM’s position as the first comprehensive system specifically designed for wireless LLM service delivery.
Appendix D illustrates the physical implementation of the WiLLM testbed, depicting the comprehensive hardware architecture that enables end-to-end LLM service delivery over wireless networks. It also showcases the WiLLM monitoring and performance analysis GUI, which provides real-time visualization and analytical capabilities for system optimization.
Appendix E details WiLLM’s hierarchical API architecture and presents the Tree-Branch-Fruit Slicing algorithm for UE resource allocation. These technical specifications elucidate the system’s cross-layer integration mechanisms and resource distribution methodologies.
Appendix F contains supplementary data analysis and visualizations from the WiLLM dataset. These figures provide deeper insights into the stability of experimental conditions, comparative performance between slice-enabled and normal traffic, and the complex interactions between various system parameters. More importantly, we conducted a causal analysis of the results shown in Figure. 6 to Figure  10 of the dataset.
Appendix G introduces two novel benchmarking metrics - the LLM-Aware Resource Efficiency Index (LAREI) and the LLM Slice Efficiency Quotient (LSEQ). These metrics are derived from rigorous analysis of the WiLLM dataset and provide a comprehensive framework for evaluating the performance of LLM wireless communication systems.
Appendix H presents a comprehensive tabulation of all metrics collected in the WiLLM dataset across the UE, RAN, and Edge Server layers. This exhaustive catalog enables researchers to understand the full scope and granularity of the empirical data foundation provided by WiLLM.
The appendix aims to provide a rigorous technical foundation for the innovative architectural designs and empirical findings presented in the main paper. By offering detailed comparative analysis, comprehensive system specifications, in-depth dataset insights, and novel performance evaluation frameworks, this supplementary material enables researchers to fully engage with the technical contributions of WiLLM and build upon its open-source foundation to advance the state of the art in wireless LLM service delivery.
Appendix B Appendix for Section 2 (Related Work and Comparison): General Comparation of Systems
Table 1. Comparison of Open Source Wireless Communication Systems and LLM deployment Systems
System Name Network Architecture Slice Support LLM Integration Resource Scheduling UE Compatibility API Framework Deployment Flexibility Dataset & Benchmark
WiLLM UE-gNB-CN with Edge Server “Tree-Branch-Fruit” Slice Designed specifically for LLM services Dual-mode resource schedule All UE Cross-layer API framework Core and edge part flexibility
OAI (Nikaein et al., 2014) UE-gNB-CN Base slice Basic resource schedule Slice-enabled UE gNB Only Traditional service-oriented
srsRAN (Flakowski et al., 2023) UE-gNB-CN Base slice Flexible resource schedule Slice-enabled UE gNB Only Traditional service-oriented
Open5GS (Open5GS, 2025) UE-gNB-CN Base slice Basic resource schedule Slice-enabled UE gNB Only Traditional service-oriented
Hugging Face’s TGI (Jain, 2022) LLM only Fine-tuning and optimization Computation Only Server
Ollama (Ollama, 2025) LLM only LLM deployment Computation Only Server
TensorRT-LLM (NVIDIA, 2025) LLM only Lightweight reasoning Computation Only Server
DeepSpeed (Rasley et al., 2020) LLM only Distributed training and reasoning Computation Only Server
Legend:  Support   Not Considered  Blue text: key comparisons
Note: This comparison contrasts systems with different foundations: OAI, srsRAN, and Open5GS are ground-up implementations, while WiLLM strategically extends OAI’s architecture. While this may appear uneven, the comparison serves to highlight functional gaps in existing systems and how WiLLM addresses these through targeted extensions. This approach enables rapid innovation while leveraging established frameworks, demonstrating how focused architectural evolution can efficiently address emerging application needs.
Table 1 presents a comprehensive comparative analysis of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks. This systematic evaluation establishes WiLLM’s distinctive contributions within the telecommunications-AI convergence landscape. While traditional systems (OAI, srsRAN, Open5GS) provide foundational wireless infrastructure components, they exhibit significant limitations in LLM integration capabilities. Concurrently, established LLM frameworks (Hugging Face’s TGI, Ollama, TensorRT-LLM, DeepSpeed) offer sophisticated model deployment but lack wireless communication integration. WiLLM uniquely bridges this interdisciplinary gap through its ”Tree-Branch-Fruit” slicing architecture, cross-layer API framework, and universal UE compatibility mechanisms. This comparative evaluation demonstrates how WiLLM transcends the limitations of existing systems through its specialized design for LLM service delivery in wireless environments.
Appendix C Appendix for Section 2 (Related Work and Comparison): Detailed Technical Comparison of Systems
Table 2. Technical Comparison of Wireless Systems
Technical Features WiLLM OAI srsRAN Open5GS FlexRIC Janus
Type LLM-Oriented Platform Open-Source 5G System Open-Source 4G/5G Suite Open-Source 5G CN Programmable RAN SDK Programmable RAN Platform
Main Features Specialized for LLM Comm. Research and Prototyping eNodeB and UE Apps LTE/NR Core Network xAPP Development O-RAN Service Model
LLM-Specific Slicing Architecture
Dynamic Slice Compatibility
Universal UE Compatibility for Slice
Multi-UE-Multi-Slice Coordination
Dual-Mode Resource Scheduling
Cross-Layer API Framework
Flexible LLM Deployment
LLM Communication Dataset
LLM Communication Benchmark
Hierarchical Slice Policy Enforcement
Application-Layer Slice Access
Synchronized Multi-Interface Metrics
Offline and Online Slice Optimization
Programmability Interface
xApp/rApp Support
Disaggregated RAN Architecture
RAN Intelligence Controller Integration
Service Model Flexibility
Multi-vendor RAN Integration
Low-latency Optimization Tools
O-RAN Compliance
Protocol Strictness
Original Protocol Stack Implementation
Original Hardware Support
Original System Architecture
Open-Source Community Activity
Standard Compliance Testing Tools
Cost Efficiency for Deployment
Legend:  Full support/Good   Not supported/Bad   Partial support/Mediocre
Note: WiLLM is an extend of OAI
Table 2 extends our comparative analysis through a granular examination of specific technical capabilities across contemporary wireless systems. This evaluation reveals significant limitations in existing platforms regarding LLM service integration. While established systems like OAI, srsRAN, and Open5GS offer partial implementation of core wireless networking features, they demonstrate substantial deficiencies in specialized LLM communication capabilities. WiLLM distinguishes itself through full implementation of LLM-specific features, including dynamic slice compatibility, multi-UE-multi-slice coordination, and hierarchical slice policy enforcement. The matrix also highlights WiLLM’s innovative application-layer slice access mechanism, which enables universal device compatibility without requiring protocol-level modifications. This detailed technical comparison substantiates WiLLM’s position as the first comprehensive system specifically designed for wireless LLM service delivery.
Appendix D Appendix for Section 4 (System): The Hardware Platform
Figure 14. Hardware deployment of the WiLLM testbed, showcasing the integrated system with core network servers, edge computing nodes, Radio Access Network equipment, and multiple user terminal devices.
Figure 14 depicts the physical implementation of our WiLLM testbed, illustrating the comprehensive hardware architecture that enables end-to-end LLM service delivery over wireless networks. The deployment integrates multiple hierarchical components: an Intel Core i9-14900KF processor serving as the gNB host, a USRP B210 software-defined radio functioning as the RF frontend, and an NVIDIA RTX 4090 GPU enabling efficient LLM inference processing. The system supports simultaneous multi-slice and multi-UE operations through specialized hardware components and SIM card integration. This hardware configuration establishes a complete signal chain from core network to edge computing infrastructure, facilitating comprehensive experimental validation of the WiLLM architecture across diverse deployment scenarios and computational requirements.
Figure 15. WiLLM monitoring and performance analysis GUI, supporting multi-dimensional metric visualization, time-series analysis, and aggregate function computation.
Figure 15 illustrates the WiLLM monitoring and performance analysis GUI, which provides real-time visualization and analytical capabilities for system optimization. The interface implements a multi-dimensional metric visualization framework supporting time-series analysis and aggregate function computation. Key functional components include: (1) communication data display for real-time performance monitoring, (2) metrics selection interface for parameter filtering and comparison, (3) time interval selection for temporal analysis, (4) aggregate function selector for statistical computation, and (5) data selection mechanisms for customizing visualization parameters. This integrated monitoring framework enables comprehensive performance evaluation across multiple system layers, facilitating both runtime optimization and longitudinal performance analysis.
Appendix E Appendix for Section 4 (System): The API Architecture and Slice Algorithm of WiLLM
Figure 16. API Architecture of WiLLM.
Figure 16 illustrates WiLLM’s hierarchical API architecture with a three-tier structure: (1) User Management API handling registration, configuration and preferences; (2) System Management API orchestrating slice operations through availability checks, request processing, and status monitoring; and (3) Resource Management API managing resource discovery, allocation and telemetry. This architecture implements cross-layer integration while maintaining separation of concerns, using RESTful APIs for synchronous operations and WebSocket connections for asynchronous monitoring.
Algorithm 1 presents the Tree-Branch-Fruit Slicing algorithm for UE resource allocation. It processes UE identity, PRBs, slice policies, channel quality indicators, and throughput metrics through four phases: (1) branch slice determination using NSSAI configuration (lines 1-4); (2) transport block size computation with proportional fair metrics (lines 5-7); (3) hierarchical policy enforcement across branches and fruits (lines 8-13); and (4) final resource allocation with MCS selection (lines 14-16). This implementation ensures efficient resource distribution while maintaining slice isolation and service differentiation across multiple UEs.
Algorithm 1 Tree–Branch–Fruit Slicing for UEs
0:  UE $u$, total PRBs $N_{\text{PRB}}$, slice policies $\mathcal{P}$, channel quality parameters $Q_{m},\,R,\,n_{\text{RB}},\,n_{\text{sym}},\,L$, and historical throughput $\Theta(u)$.
0:  Final allocated PRB count $R(u)$ and selected MCS for $u$.
1:  /* Determine UE’s branch slice based on NSSAI configuration */
2:  Define the slice set of $u$ as
$\mathcal{S}(u)=\{\,s\mid s\text{ is extracted from }\mathrm{NSSAI}(u)\,\}$
3:  Compute the branch slice index
$b=\operatorname{MatchBranch}(\mathcal{S}(u),\mathcal{P})$
4:  Retrieve branch policy parameters:
$\alpha_{b}^{\min}=\mathcal{P}(b).\min\_ratio\quad\&\quad\alpha_{b}^{\max}=\mathcal{P}(b).\max\_ratio$
5:  /* Compute Transport Block Size using channel parameters */
$\text{TBS}(u)=f\bigl{(}Q_{m},\,R,\,n_{\text{RB}},\,n_{\text{sym}},\,L\bigr{)}$
6:  Compute the proportional fair metric:
$\gamma(u)=\frac{\text{TBS}(u)}{\Theta(u)}$
7:  Derive the initial allocation proportion via a function $\phi(\cdot)$:
$r_{\text{init}}=N_{\text{PRB}}\cdot\phi\bigl{(}\gamma(u)\bigr{)}$
8:  Enforce branch policy constraints (lower and upper):
$r_{\text{branch}}=\min\Bigl{\{}\,r_{\text{init}},\;N_{\text{PRB}}\cdot\alpha_{b}^{\max}\Bigr{\}}$
$r_{\text{branch}}=\max\Bigl{\{}\,r_{\text{branch}},\;N_{\text{PRB}}\cdot\alpha_{b}^{\min}\Bigr{\}}$
9:  if a fruit mapping exists for $u$ then
10:     Let $(\pi(u),\;r_{\min},\;r_{\max})$ be the fruit parameters for $u$.
11:  else
12:     Define default fruit parameters:
$\pi(u)=1,\quad r_{\min}=N_{\text{PRB}}\cdot\alpha_{b}^{\min},\quad r_{\max}=N_{\text{PRB}}\cdot\alpha_{b}^{\max}$
13:  end if
14:  Final resource allocation:
$R(u)=\min\Bigl{\{}\,\max\Bigl{\{}\,\pi(u)\cdot r_{\text{branch}},\;r_{\min}\Bigr{\}},\;r_{\max}\Bigr{\}}$
15:  Determine MCS from channel conditions:
$\text{MCS}=\operatorname{SelectMCS}(u,Q_{m},R,L)$
16:  return  $(R(u),\,\text{MCS})$
Appendix F Appendix for Section 5 (Dataset)
Figure 17. The variations in SNR during the data collection process.
Figure 18. The variations in PH_dB during the data collection process.
Figure 19. The variations in UL_THR during the data collection process.
Figures 17 and  18 present temporal analyses of Signal-to-Noise Ratio (SNR) and Power Headroom (PH_dB) variations during the data collection process, respectively. These visualizations demonstrate the stability of the experimental environment across multiple measurement dimensions. Figure 17 reveals consistent SNR levels (mean: 150.76, standard deviation: 21.61) across the data collection period, with minimal transient fluctuations, confirming reliable signal quality for experimental validity. Similarly, Figure 18 shows highly stable power headroom measurements (mean: 59.41, standard deviation: 2.41), indicating consistent UE transmission capabilities throughout the experimental process. These stability metrics validate the integrity of our dataset by confirming that the observed performance variations result from intentional experimental manipulations rather than environmental fluctuations or hardware inconsistencies.
Figure 19 displays comparative analyses of uplink throughput between normal traffic (top) and slice-enabled traffic (bottom) during the data collection process. The slice-enabled implementation demonstrates significantly higher average throughput compared to normal traffic, representing a 43.5% performance improvement. The slice-enabled approach also shows more dynamic adaptation with higher peak throughput, enabling better responsiveness to varying service demands. While normal traffic exhibits regular but constrained patterns, the slice-enabled traffic shows more variable but overall superior performance characteristics with distinct periodic bursts that effectively utilize available bandwidth. This empirical evidence confirms that WiLLM’s slicing architecture successfully enhances throughput efficiency, allowing more responsive performance scaling for LLM services while maintaining an appropriate baseline throughput level, which is crucial for consistent user experience in latency-sensitive LLM applications.
F.1. Analysis of Uplink vs. Downlink Transmission Patterns (Figures 6, 7 and 8)
The differential performance profiles across slice configurations illustrated in Figure 8 substantiate the efficacy of WiLLM’s “Tree-Branch-Fruit” slicing architecture, with differences primarily attributable to input modality variations. The pronounced disparity in inference time contribution (11.6% to 59.6%) stems fundamentally from the computational asymmetry between processing graphical inputs in uplink scenarios versus textual inputs in downlink scenarios. Image-based queries in uplink transmissions necessitate substantially higher computational overhead for visual feature extraction, semantic interpretation, and multimodal reasoning, consequently dominating the latency profile. Conversely, text-based inputs in downlink scenarios require comparatively less computational intensity, shifting the performance bottleneck toward transmission components (28.9% to 54.7% variation in transmission time). This input-dependent computational variance underscores the necessity of WiLLM’s specialized slicing approach, which adaptively allocates resources based on modality-specific requirements rather than uniform distribution models. The statistically significant performance differentials between slices confirm that WiLLM effectively maintains service isolation while accommodating the inherent computational asymmetry between visual and textual processing pipelines, validating the fundamental premise of modality-optimized “Fruit Slices” within the hierarchical resource management framework.
It is important to note that our current implementation does not utilize specialized image generation models, which would likely introduce additional computational demands in the downlink scenario. Such models might shift the bottleneck toward inference processing rather than transmission, potentially altering the observed latency distribution patterns. Nevertheless, the empirical evidence clearly demonstrates that downlink image transmission presents a more significant bandwidth challenge than uplink image transmission, likely due to differences in compression efficiency and quality requirements between input and output images.
F.2. Resource Allocation-Utilization Relationship Analysis (Figures 9 and 10)
The non-linear correlation between PRB allocation patterns (Figure 9) and actual byte transmission volumes (Figure 10) reveals critical insights into resource utilization efficiency within wireless LLM service delivery. While WiLLM successfully implements differentiated PRB allocation conforming to slice policies, the observed discrepancy between allocated resources and realized throughput reflects complex cross-layer interdependencies spanning the entire system architecture. This phenomenon stems from multi-dimensional factors: (1) variable UE request patterns and buffer state fluctuations that create unpredictable traffic bursts; (2) server-side processing variability due to LLM inference time inconsistencies and computational resource contention; and (3) system execution mechanisms within the gNB layer, including scheduler implementation details, slice prioritization algorithms, and HARQ management policies. The visualization of slice-distinguished traffic with periodic transmission bursts demonstrates how these factors propagate across architectural boundaries, creating emergent system behaviors that cannot be predicted through isolated layer analysis. This empirical evidence substantiates the necessity of WiLLM’s cross-layer API framework, which enables coordinated adaptation across UE configuration, gNB resource allocation, and server processing prioritization based on observed end-to-end performance characteristics rather than theoretical layer-specific models—establishing a comprehensive feedback system essential for optimizing LLM service delivery in wireless environments.
F.3. Dataset Collection Parameters
F.3.1. UE Configuration Parameters
During data collection, we employed a configurable UE client with the following parameter ranges:
Parameter Description Values
Image Capture Resolution Resolution of images captured by smart glasses 320$\times$240 to 640$\times$480 pixels
Display Resolution Resolution of the glasses display panel 800$\times$600 to 1280$\times$720 pixels
Request Mode Type of request sent to LLM “image_request”, “text_request”
LLM Model Model used for inference “llava”, “llama3.2”
Response Word Count Maximum words requested in LLM response 50, 100, 150, 200 words
Request Frequency Periodic request interval 5 seconds (default)
Table 3. UE configuration parameters used in dataset collection
Our experiment controller dynamically adjusted these parameters with varying coefficients (1.0, 0.9, 0.8, 0.7, 0.6, 0.5) applied to the base resolutions to analyze performance across different capture qualities. This approach simulated diverse real-world usage scenarios with varying resource constraints and quality requirements.
F.3.2. Network Slicing Configuration
The network slicing configuration used in our experiments consisted of three primary slices with different resource allocation parameters. Three fruit slices were defined with max_ratio values of $\{30\%,60\%,90\%\}$, all associated with the first parent slice. During data collection, our slice controller cycled through these configurations at 30-second intervals, updating the slice mapping for UEs to create varied resource allocation patterns throughout the dataset collection process.
Appendix G Benchmarking Metrics for LLM Wireless Communication
Figure 20. Latency, scheduled uplink bytes, and request bytes correlation heatmap. This figure presents a heatmap visualization of the complex interdependencies between latency, scheduled uplink bytes, and request bytes. The color-coded matrix reveals non-linear relationships between these parameters, identifying potential optimization opportunities and constraint boundaries that inform the development of comprehensive performance metrics for LLM wireless systems.
Figure 21. Correlation heatmap between slice configuration, request bytes, and uplink BLER. This figure presents a multi-dimensional analysis of how slice configuration, request payload size, and resulting error rates interact. The color-coded visualization provides empirical evidence for non-trivial relationships between these parameters, supporting the development of a slice efficiency index that balances resource allocation against transmission reliability.
G.1. LLM-Aware Resource Efficiency Index
Based on Figure.21, our dataset analysis reveals complex interdependencies between latency, scheduled uplink bytes, and request bytes. The visualization demonstrates non-linear relationships between these critical parameters, suggesting the existence of multi-dimensional optimization boundaries that traditional isolated metrics fail to capture. The color-coded matrix illuminates several key observations:
  1. (1)
    Regions of optimal performance emerge at specific combinations of resource allocation and request volume, rather than following a monotonic improvement pattern with increasing resources.
  2. (2)
    Latency exhibits plateau effects at certain throughput thresholds, indicating saturation points where additional resources yield diminishing returns.
  3. (3)
    Request volume influences the relationship between resource allocation and latency in non-trivial ways, creating complex performance landscapes that require holistic optimization.
It is important to note that our heatmap representation, constrained by dimensionality limitations of visualization, necessarily focuses on three representative parameters. These selected metrics—latency, scheduled uplink bytes, and request bytes—serve as proxies for more complex system characteristics. For instance, scheduled uplink bytes approximate but do not completely capture overall resource utilization patterns across the network stack, while request bytes provide insight into workload demands without fully representing request resolution or complexity variations. Nevertheless, these parameters offer meaningful approximations of the fundamental system dynamics at play.
To quantify these relationships, we propose the LLM-Aware Resource Efficiency Index (LAREI), formulated as:
$\text{LAREI}=\frac{\text{RDV}\cdot\log(1+\text{LLM\_Para})}{\text{Resources}\cdot\text{Latency}}\cdot\omega$
Where:
  • RDV (Request Data Volume) quantifies the amount of data requested by the user (approximated by request bytes in our dataset)
  • LLM_Para (LLM Parameter Count) represents the computational complexity of the language model (measured in billions of parameters)
  • Resources (Allocated Communication Resources) represents the network resources provisioned for transmission (approximated by scheduled uplink bytes)
  • Latency measures end-to-end response time
  • $\omega$ is a normalization coefficient that adjusts for system-specific characteristics
This formulation captures the fundamental efficiency relationship observed in our dataset while incorporating the computational dimension of LLM services. The logarithmic scaling of the parameter count acknowledges the empirically observed sub-linear relationship between model size and perceived quality improvements. This multidimensional index effectively balances the communication-computation trade-off inherent in LLM service delivery over wireless networks.
The LAREI serves as a comprehensive metric for evaluating resource allocation algorithms specifically tailored to LLM communication services, enabling researchers to identify optimal operating configurations that balance communication efficiency with model sophistication.
G.2. LLM Slice Efficiency Quotient
Figure. 9, Figure 10 and Figure. 21 reveal critical insights into network slicing performance that challenge conventional resource allocation assumptions. It demonstrates that higher resource allocation does not necessarily translate to proportionally lower error rates, suggesting an efficiency frontier in slice optimization. The comparative visualization of BLER between slice-enabled and normal traffic illustrates that excessive resource provisioning can potentially increase transmission errors.
Figure. 21 further illuminates this phenomenon through a correlation heatmap between slice configuration, request bytes, and uplink BLER. The visualization reveals distinct patterns where certain slice configurations exhibit superior error performance despite lower resource allocation, while others show degraded performance despite abundant resources. This empirical evidence supports the development of a slice efficiency metric that transcends simple resource allocation quantities.
It should be emphasized that our heatmap analysis, while instructive, represents a dimensional reduction of a multifaceted system. The three parameters visualized—slice configuration, request bytes, and BLER—function as indicative proxies for the broader system dynamics. For example, slice configuration encapsulates numerous implementation details not fully represented in the visualization, while BLER provides insight into transmission reliability without comprehensively capturing all aspects of service quality. Despite these limitations, the observed patterns reveal fundamental relationships that inform our metric development.
Based on these observations, we propose the LLM Slice Efficiency Quotient (LSEQ):
LSEQ=RDV_slice(1Error Rate)LLM_Para_sliceSlice $\text{LSEQ}=\frac{\text{RDV\_slice}\cdot(1-\text{Error Rate})\cdot\sqrt{\text{LLM\_Para\_slice}}}{\text{Slice Resources}}\cdot\delta$
Where:
  • RDV_slice represents the amount of data requested by users within the slice
  • Error Rate quantifies transmission errors (represented by BLER in our dataset)
  • LLM_Para_slice reflects the computational complexity of the language model deployed within the slice
  • Slice Resources measures the communication resources provisioned to the slice
  • $\delta$ is a calibration parameter that normalizes across different deployment scenarios
[…]
The LSEQ enables researchers to evaluate slicing algorithms specifically designed for LLM service delivery, accounting for the complex interplay between resource allocation, transmission reliability, and model sophistication. It provides a theoretically sound framework for optimizing slice configurations in LLM-oriented wireless communication systems.
Appendix H All Matrics of Dataset
Table 4. UE Layer Dataset Metrics
Layer Metric Description
UE Timestamp - Request initiation timestamp (Unix epoch milliseconds)
Wireless Comm Time - UE-gNB air interface duration (PDCP layer measurement)
Total Comm Time - UE-side end-to-end latency (request to response)
Tx Image Resolution - Original image dimensions (e.g., 1920×1080)
Rx Image Resolution - Received image dimensions post-processing
Expected Word Count - User-specified desired explanation length
Actual Word Count - LLM-generated response word count
LLM Model - Model architecture (LLaMA, LLaVA, etc.)
Request Mode - image_request or text_request
Upload Periodicity - Uplink interval (ms, 0=event-driven)
Uplink Time - UE-to-network latency (RLC layer)
Downlink Time - Network-to-UE latency (PDCP layer)
Downlink Text Size - Response payload size (UTF-8 bytes)
Uplink Bytes - Total bytes of the request (uplink)
Downlink Bytes - Total bytes of the response (downlink)
Table 5. Edge Server Layer Dataset Metrics
Layer Metric Description
Server LLM Inference Time - Model forward pass latency (ms)
Server Processing Time - Processing time in server
Input Tokens - Tokens consumed by LLM
Output Tokens - Tokens generated by LLM
Cold Start Time - Disk-to-RAM loading time
Warm Start Time - Cache reload time
BLEU Score - Text quality evaluation metric
ROUGE Score - Recall-based text evaluation
Semantic Score - Embedding similarity score
GPU Utilization - GPU compute usage (%)
VRAM Usage - Video memory consumption
Downlink Image - Base64 output image
Response Text - LLM-generated response
Table 6. RAN Layer Dataset Metrics
Layer Metric Description
RAN gNB Timestamp - gNB timestamp alignment (Unix epoch ms)
Frame Number - 5G NR system frame (0-1023)
Slot Number - Within-frame slot index (0-159)
IMSI - International Mobile Subscriber Identity
RNTI - Radio Network Temporary Identifier
UE ID - Device identifier within gNB
UE Number - Logical connection index
DL Throughput - Instantaneous downlink rate (Mbps)
UL Throughput - Instantaneous uplink rate (Mbps)
PH (dB) - UE Power Headroom report
PCMAX (dBm) - Max transmission power
Avg RSRP - Reference Signal Received Power
CQI - Channel Quality Indicator (0-15)
RI - MIMO Rank Indicator
DL MCS - Downlink Modulation and Coding Scheme
UL MCS - Uplink Modulation and Coding Scheme
Scheduled UL Bytes - Uplink buffer status
Estimated UL Buffer - gNB buffer estimation (bytes)
DL PDUs Total - Aggregated downlink PDUs
DL BLER - Downlink Block Error Rate (%)
UL BLER - Uplink Block Error Rate (%)
DL-SCH Bytes - Downlink shared channel payload
DL-SCH RBs - Downlink Resource Blocks allocated
UL-SCH Bytes - Uplink shared channel payload
UL-SCH RBs - Uplink Resource Blocks allocated
UL MAC SDUs - MAC Service Data Units count
Primary Slice Max - Main slice maximum resources
Primary Slice Min - Main slice minimum guarantee
Secondary Slice Max - Secondary slice maximum
Secondary Slice Min - Secondary slice minimum

### v2

@@
-WiLLM: An Open Wireless LLM Communication System
+WiLLM: an Open Framework for LLM Services over Wireless Systems
 Abstract.
-The rapid evolution of Large Language Models (LLMs) threatens to overwhelm the existing wireless infrastructure, which calls for architectural innovations and frameworks to support the surging mobile LLM services.
-This paper proposes WiLLM, the first open-source wireless communication system specifically designed for LLM services, consisting of three progressive innovations.
-Firstly, we establish a novel paradigm by deploying LLMs in core networks (CNs) with abundant GPUs to enable distributed LLM inference services, strategically positioning LLM inference at the convergence point of the backbone bandwidth and the edges of the cellular network.
-Secondly, we propose an innovative extension to the conventional network slicing architecture, redefining its design with a “Tree-Branch-Fruit” framework.
-Its specialized adaptation for LLM services enables telecom operators to monetize these services through slice subscriptions while retaining infrastructure ownership.
-Finally, to operationalize this architectural vision, we present WiLLM, which systematically addresses critical limitations in existing solutions through novel key capabilities previously unattainable: (1) enhanced slice orchestration, extending existing primary slicing to implement a dual-layer slicing architecture and achieving coordinated multi-UE-multi-slice scheduling mechanisms that provide finer-grained resource allocation for LLM services; (2) universal UE compatibility through application-layer tunneling mechanisms, enabling legacy devices without native slicing capabilities to utilize LLM slice services without hardware upgrades; (3) dual-mode scheduling and cross-layer APIs, supporting flexible deployment from CNs to servers.
-Built on OpenAirInterface, WiLLM extends this established framework with innovations specific to LLM services, lowering the barrier for adoption by researchers without wireless expertise.
-Furthermore, we release the first LLM wireless communication dataset with 1,649,996 records and synchronized 58-dimensional metrics, along with two relevant benchmarks.
-A small but quintessential case study of smart glasses demonstrates practical viability in resource-constrained devices and human-machine interaction scenarios.
-WiLLM aims to build an open platform and ecosystem for cross-layer optimization and AI-telecom convergence.
-The code, datasets, hardwares and videos are available at https://openwillm.github.io.
+Large Language Model (LLM) services fundamentally differ from traditional Deep Neural Network (DNN) applications in wireless networks.
+We identify three critical distinctions: (1) unlike traditional DNNs with unidirectional data flows, LLM’s multimodal interactions create bidirectional heavy loads with contrasting bottlenecks, requiring direction-aware resource scheduling; (2) while traditional DNNs exhibit fixed computational patterns, LLM’s highly variable inference times interact complexly with network slicing, causing dynamic bottleneck migration; and (3) in contrast to predictable DNN traffic, LLM’s token streams demonstrate unprecedented burstiness and state dependencies.
+These insights motivate WiLLM, the first open-source framework, implemented as a wireless platform, for LLM service research.
+Built on OpenAirInterface, WiLLM introduces several technical innovations: dynamic slice compatibility, universal UE compatibility through application-layer tunneling, multi-UE multi-slice scheduling, dual-mode resource allocation, and cross-layer APIs.
+In addition, WiLLM eliminates the need for specialized wireless expertise, enabling researchers and developers to experiment with LLM services over realistic cellular networks.
+We demonstrate the platform’s capabilities through a smart glasses case study and provide a comprehensive dataset of 1̃.6 million synchronized measurements.
+The complete system, dataset, and appendix are available at https://openwillm.github.io.
 Large Language Model, Wireless Communication System
 copyright: none
+1. Introduction
+The proliferation of Large Language Models across mobile applications creates unprecedented demands on wireless network infrastructure.
+From augmented reality assistants to real-time translation services, LLM-powered applications require seamless integration of computational inference and wireless communication (Bariah et al., 2024).
+However, deploying LLM services over cellular networks reveals fundamental mismatches between their operational characteristics and existing network architectures (Boateng et al., 2024; Chaoub and Elkotob, 2025).
+Modern wireless systems, evolved to handle video streaming and web traffic, struggle with the unique patterns of token generation, multimodal processing, and bidirectional heavy loads that characterize LLM interactions.
 Figure 1. Contributions and advantages of the proposed WiLLM.
-1. Introduction
-The convergence of artificial intelligence and wireless communications represents one of the most significant technological intersections of the current decade.
-As computational capabilities and algorithmic sophistication continue to advance, LLMs have emerged as transformative tools with far-reaching implications across diverse application domains.
-The integration of these models into wireless communication systems has catalyzed a paradigm shift, enabling advanced intelligent services ranging from real-time multilingual translation to context-aware augmented reality applications (Bariah et al., 2024).
-As this integration accelerates, telecommunication infrastructure faces unprecedented challenges in supporting the computational demands of these resource-intensive models.
-Current deployments predominantly rely on either edge computing paradigms or centralized cloud processing approaches, each presenting significant limitations.
-Edge deployment minimizes transmission latency but suffers from constrained computational capacity and thermal limitations in user equipment (UE), while cloud-based solutions introduce unpredictable internet latency that severely impacts quality-of-experience (QoE) for interactive LLM services (Zheng et al., 2024).
-Moreover, the exponential growth in concurrent LLM service requests threatens to overwhelm both communication channels and computational resources, creating critical bottlenecks in service provisioning.
-Recent implementations across several institutions in China, including deployments of Deepseek at multiple universities (Engineer, 2025), demonstrate the feasibility of local LLM deployments at the network edge (Boateng et al., 2024).
-While these edge deployments represent a promising step toward LLM integration within telecommunications infrastructure, they still encounter resource limitations and scalability constraints inherent to edge computing.
-These implementations nonetheless validate the technical viability of localized LLM services and establish a foundation for our proposed migration of LLM computational resources to the core network, where greater processing capacity and more efficient resource allocation can be leveraged.
-This transition from edge to core network deployment represents a natural evolution in the LLM service architecture, addressing the scalability limitations of edge deployments while maintaining the latency advantages over cloud-based solutions (Chaoub and Elkotob, 2025).
-The current network architecture exhibits three fundamental drawbacks hindering effective LLM service integration: (1) the absence of dedicated open testbeds for LLM-specific communication scenarios, (2) the inherent conflict between centralized computation paradigms and distributed network architectures, and (3) the lack of viable business models for telecom operators in LLM service provisioning.
-Existing network slicing architectures provide potential foundations for addressing these challenges through resource isolation and service customization.
-However, conventional slicing methodologies primarily focus on traditional communication services and lack the specialized mechanisms required for computationally intensive LLM workloads (Dehkordi et al., 2023).
-Furthermore, the absence of standardized APIs and resource management frameworks inhibits the development of cross-layer optimization strategies essential for efficient LLM service delivery.
-Another critical limitation is the high usability barrier of current systems.
-Existing platforms require specialized wireless expertise and native slicing support on user equipment, restricting access primarily to communication specialists (Zheng et al., 2024).
-This complexity explains why testbeds are often used only by their developer communities, significantly limiting cross-disciplinary research and broader adoption for LLM service experimentation.
-Moreover, the field lacks comprehensive empirical datasets capturing cross-layer performance metrics, standardized benchmarking methodologies for systematic comparison, and representative case studies demonstrating viability in resource-constrained environments.
-This empirical deficit perpetuates the disconnect between algorithmic innovation and operational deployment, hampering the development of generalizable principles for wireless LLM communication systems.
-Motivated by the above system and networking challenges, we present WiLLM, the world’s first open-source wireless communication system specifically designed for mobile LLM services.
-Building upon the solid foundation provided by OpenAirInterface’s modular architecture, WiLLM strategically extends and enhances this established framework with specialized components for LLM service integration, creating a system that maintains compatibility with existing infrastructure while introducing transformative capabilities for AI-driven communications research.
-The work unifies the following progressive contributions:
-1.1. Paradigm
-Novel Network Paradigm for LLM Services (Section 3).
-WiLLM establishes a novel paradigm by deploying LLMs in core networks (CNs) with GPU infrastructure, enabling distributed LLM inference services strategically positioned at the convergence point between high-bandwidth backbone connections and cellular network edges.
-This architectural configuration optimizes computational-communication tradeoffs while facilitating operator monetization through subscription-based network slice allocation.
-1.2. Architecture
-“Tree-Branch-Fruit” Slicing Architecture (Section 3.3).
-We propose a multi-layered slicing architecture wherein the “tree” represents foundational network infrastructure, the “branches” correspond to conventional communication service slices, and the “fruits” denote specialized LLM service slices.
-This architecture establishes a “slice-as-LLM-service” paradigm while maintaining compatibility with existing slicing mechanisms, achieving coordinated scheduling across multiple users and slices through cross-layer resource optimization.
-1.3. System
-The Wireless Communication System Specialized for LLM Service (Section 4).
-Based on OpenAirInterface, WiLLM implements five pivotal innovations: dynamic slice compatibility, universal UE compatibility, dual-mode resource scheduling, cross-layer API framework, and flexible deployment support.
-This integrated solution addresses critical limitations in existing wireless infrastructure while preserving backward compatibility.
-1.4. Dataset
-The World’s First LLM Wireless Communication Dataset and Benchmarks (Section 5).
-We introduce the first comprehensive LLM communication dataset with 1,649,996 records, each containing 58 communication metrics synchronized across UE-gNB-CN interfaces.
-This enables holistic cross-layer analysis of LLM communication patterns.
-In addition, we present two metrics, LAREI and LSEQ, as benchmarks that offer a standardized methodology for wireless LLM service evaluation.
-1.5. Case Study
-Smart Glasses Case Study (Section 6).
-The practical utility of WiLLM is validated through a smart glasses application that implements gesture-triggered queries and scene interpretation.
-Based on the principle of user-perceived continuity in human-computer interaction, we establish a response latency target of approximately 2 seconds, achieving optimal slicing through offline statistical analysis and online learning methodologies.
-2. Related Work and Comparison
-2.1. Wireless Communication for LLM
-The convergence of LLMs with wireless systems has precipitated prolific theoretical advancements characterized by a critical implementation deficit.
-Despite many approaches being proposed, the field confronts a fundamental disjunction between algorithmic innovation and operational realization.
-Architectural paradigms manifest in three configurations: 1.
-Cloud-edge frameworks (Chen et al., 2024b) (Wu et al., 2020) (Dong et al., 2024) establish a hierarchical computational distribution by allocating processing tasks between the cloud and edge devices.
-This approach combines the high computational power of the cloud with the low latency characteristics of the edge.
-For example, NetGPT (Chen et al., 2024b) relies on cloud-edge collaboration in personalized generative services.
-However, the dynamic nature of wireless channels limits its adaptability.
-Reference (Wu et al., 2020) explores the potential of cloud-edge computing, while (Dong et al., 2024) highlights the challenges of implementing edge AI generated from cloud-based LLMs.
-2.
-Edge computing deployment (Cai et al., 2024) (Nelson and Toloui, 2023) (Tang et al., 2023) runs LLMs directly on edge servers or devices, enhancing performance through quantization techniques.
-For instance, Edge-LLM (Cai et al., 2024) optimizes edge adaptability using unified compression and adaptive layer voting, while (Nelson and Toloui, 2023) demonstrates hardware-supported edge deployment cases.
-Nevertheless, the limited capacity of edge devices remains a bottleneck, as further validated in (Tang et al., 2023).
-3.
-Distributed architectures (Xue et al., 2024) (Mitra et al., 2024) decompose LLM computational tasks across multiple nodes to fully utilize distributed resources.
-For example, WDMoE (Xue et al., 2024) proposes a wireless distributed framework based on mixture-of-experts, though its validation in congested scenarios remains insufficient.
-Resource optimization methodologies span cross-modal approaches (JPPO (You et al., 2024)) integrating prompt compression with power optimization; adaptive partitioning strategies (Chen et al., 2024a) reconfiguring computational boundaries under network fluctuations; and collaborative frameworks (He et al., 2024) optimizing offloading decisions across distributed resources.
-Service integration architectures encompass specialized slicing methodologies (LLM-Slice (Liu et al., 2024)), intelligent orchestration approaches (Dandoush et al., 2024), and long-context integration strategies (Xu et al., 2025; Zhang et al., 2025).
-Bottlenecks and our solutions: The field faces three bottlenecks: theory-centric validation lacking real-world relevance, insufficient cross-layer integration, and unproven business viability.
-WiLLM addresses these through an empirical platform bridging theory and implementation.
-Unlike conventional edge server strategies that incur coordination inefficiencies, we deploy LLMs on GPU-equipped infrastructure at the backhaul-RAN nexus, enhancing computational capacity, resource optimization, and infrastructure integration.
-2.2. Related Systems and Comparisons
-In this section, we evaluate prominent open-source systems for their capabilities across a range of technical features essential for modern wireless networks, including OAI, srsRAN, Open5GS, FlexRIC, and Janus.
-While these systems contribute significantly to wireless network research and development, they exhibit limitations in fully supporting the comprehensive demands of next-generation networks.
-OAI provides a complete 4G/5G protocol stack implementation, spanning from the physical layer to the network layer.
-Its modular design supports academic research and prototyping (Nikaein et al., 2014).
-srsRAN is a lightweight and highly customizable 4G/5G software suite (Gomez-Miguelez et al., 2016), offering flexibility in scheduling and QoS management.
-Open5GS implements comprehensive 5G core network functions, including AMF, SMF, and UPF components (Barbosa et al., 2024), establishing a solid foundation for network integration.
-FlexRIC is a programmable RAN SDK focused on 5G multi-tenancy and low-latency optimization (Schmidt et al., 2021).
-Janus is a programmable RAN platform developed by Microsoft, based on the O-RAN architecture, supporting dynamic service models (Foukas et al., 2023).
-While these systems offer essential assistance in certain aspects of wireless networking, they exhibit inadequate support for general technical features like UE compatibility, slice flexibility, and specialized support for LLM communications, etc.
-Detailed objective metric comparisons are presented in Table 2 and Table 1 in the Appendix.
-2.2.1. Research Gaps and Positioning of WiLLM
-While existing open-source communication platforms provide valuable research infrastructure, comprehensive analysis reveals significant operability impediments that restrict their adoption beyond specialized research communities.
-As quantified in Table  1 and 2, current systems demonstrate substantial limitations in critical usability dimensions.
-The requirement for native slicing support at the UE level presents a particularly problematic constraint, effectively excluding standard commercial devices without specialized protocol modifications (Saboorian and Xiang, 2017; Andrade and Wickboldt, 2025).
-This limitation fundamentally contradicts the purpose of experimental testbeds, which should facilitate algorithm validation across diverse implementation scenarios rather than constraining experimental configurations to narrow protocol-specific deployments.
-Furthermore, these resulted in an undesirable ecosystem fragmentation where testbeds remain utilized primarily by their own development communities, significantly limiting cross-disciplinary innovation.
-WiLLM aims to bridge the technological gap between wireless communication and LLM systems by extending the OAI architecture to build the first open-source wireless communication system designed specifically for LLM services.
-3. Paradigm
+Figure 2. The LLM Services Pipeline over WiLLM (Liu et al., 2024).
+The challenge extends beyond simple capacity constraints.
+Our measurements across diverse deployment scenarios reveal that LLM services violate core assumptions underlying current wireless system designs.
+Traditional traffic engineering approaches assume predictable patterns, unidirectional dominance, and static computational requirements (Andrade and Wickboldt, 2025).
+LLM services break these assumptions through their generative nature, creating workloads that dynamically shift between computation and communication bottlenecks within single sessions.
+This fundamental mismatch prevents effective service delivery and optimization using existing tools and methodologies.
+Compounding these technical challenges, the research ecosystem lacks accessible platforms for studying LLM behavior in wireless environments.
+Current open source testbeds such as OpenAirInterface (Nikaein et al., 2014) and srsRAN (Gomez-Miguelez et al., 2016) require deep telecommunications expertise and specialized hardware, effectively excluding general researchers and developers from wireless experimentation.
+Meanwhile, LLM deployment frameworks like Ollama (Ollama, 2025) and TensorRT-LLM (NVIDIA, 2025) operate without awareness of wireless constraints.
+This divide between communities impedes innovation at their intersection, where the most impactful advances will emerge in mobile AI services (Zheng et al., 2024).
+Our extensive empirical analysis, detailed in Section 2, identifies specific characteristics that distinguish LLM services from traditional applications and motivate specialized system design.
+These findings drive WiLLM, an open-source platform that bridges the gap between wireless infrastructure and LLM service.
+Rather than treating LLM services as generic data traffic, WiLLM incorporates architectural innovations specifically addressing their unique requirements while maintaining accessibility for researchers without wireless expertise.
+Figure 1 and 2 illustrates WiLLM’s advantages and the service pipeline.
+Built on OpenAirInterface’s robust protocol implementation, WiLLM extends the base platform with LLM-aware capabilities while preserving compatibility with standard devices and networks.
+The system enables researchers to experiment with real cellular networks using commercial hardware, deploy various LLM models from Llama (ggml-org, 2025) to LLaVA (Liu et al., 2023), and observe detailed cross-layer interactions through comprehensive monitoring interfaces.
+These capabilities transform wireless LLM research from a specialized domain requiring extensive telecommunications knowledge to an accessible field where AI researchers can directly contribute innovations.
+To evaluate the platform’s capability, we conducted a case study with smart glasses, where a team with minimal wireless expertise successfully deployed and optimized a multimodal LLM service.
+This validation confirms that WiLLM achieves its dual objectives: providing the technical capabilities necessary for LLM service delivery while removing barriers that have historically separated wireless and AI research communities.
+This paper makes three primary contributions.
+(1) We provide empirical evidence establishing how LLM services differ fundamentally from traditional network applications, based on analysis of over 1.6 million synchronized measurements across network layers.
+(2) We present WiLLM as an open experimental platform that democratizes wireless LLM research through architectural innovations and intuitive interfaces.
+(3) We release our complete implementation, comprehensive dataset, and benchmarking tools to foster an ecosystem for wireless LLM innovation.
+The remainder of this paper is organized as follows.
+Section 2 analyzes the unique characteristics of LLM services through empirical measurements.
+Section 3 presents WiLLM’s system architecture and design principles.
+Section 4 details our implementation on OpenAirInterface.
+Section 5 validates our design through comprehensive evaluation.
+Section 6 reviews related work, and Section 7 concludes.
+2. LLM Services: Why Different from Traditional DNN over Wireless Systems?
+We deployed WiLLM to systematically characterize LLM service behavior across commercial UE devices, a 5G RAN with configurable resource scheduling, and a GPU-accelerated Core Network.
+Our experimental design combines static and dynamic configurations for both UE and gNB components across four operational scenarios, isolating the effects of system adaptability and network slicing on service performance.
+Multimodal workloads include image-to-text queries that stress uplink capacity, and text-to-image generation that creates diverse downlink requirements.
+NTP-based synchronization achieves millisecond-level accuracy for cross-layer correlation, enabling us to trace requests from application initiation through network transmission to inference completion.
+Through careful control of network conditions and computational isolation, we collected extensive empirical evidence revealing how LLM services fundamentally differ from traditional wireless applications, as detailed in the following three insights.
+2.1. Insight 1: Bidirectional Heavy Load Asymmetry
+Traditional DNN applications in wireless networks exhibit clear unidirectional patterns.
+For example, image classification services upload raw images and receive lightweight labels, while video analytics applications stream uplink data for cloud processing.
+This unidirectional dominance enables straightforward optimization: allocate resources primarily for the dominant direction.
+LLM services fundamentally break this pattern due to their multimodal interaction capabilities.
+Within a single session, users seamlessly transition between uploading images for scene understanding and requesting image generation from text descriptions.
+Our measurements reveal that this creates two contrasting operational modes with opposite bottlenecks.
+As shown in Figure 3, in uplink-heavy scenarios (image-to-text), computational inference at the LLM dominates the end-to-end latency, accounting for 74% to 87% of total delay, while network transmission contributes marginally.
+Conversely, Figure 4 demonstrates that in downlink-heavy scenarios (text-to-image), network transmission becomes the primary constraint, consuming 81% to 86% of total latency, with inference time reduced from 12% to 17%.
+This asymmetry stems from the fundamental nature of multimodal LLMs.
+Processing high-resolution images requires extensive computation for feature extraction and semantic understanding, creating uplink inference bottlenecks.
+Meanwhile, generating detailed visual content produces large data payloads that stress downlink capacity.
+The continuous nature of LLM interactions means both patterns occur within the same session, making traditional unidirectional optimization strategies ineffective.
+Systems must dynamically recognize and adapt to these contrasting requirements, a capability absent in current wireless architectures designed for predictable traffic patterns.
+2.2. Insight 2: Computation-Communication Coupling in Slicing
+Traditional DNN workloads demonstrate predictable computational patterns.
+A ResNet-50 (Kaiming He and Sun, 2015) model requires approximately 4 GFLOPs per image regardless of network conditions.
+A YOLO (Redmon et al., 2016) detector maintains consistent inference time across deployments.
+This computational predictability decouples inference from network resource allocation, enabling independent optimization of computation and communication resources.
+Figure 3.
+Component distribution and latency analysis for uplink transmissions across resolution groups.
+The x-axis labels R1-R6 represent various resolution groups for image requests sent to LLM, where R1-R6 correspond to increasing resolutions.
+The left panel shows the latency composition of different components (uplink, downlink, inference), the center panel presents statistical trends with standard errors, and the right panel quantifies each component’s percentage contribution to total latency.
+Figure 4.
+Component distribution and latency analysis for downlink transmissions across resolution groups.
+Similar to Figure 3, it provides three complementary views: component distribution, statistical latency metrics, and proportional component analysis for downlink scenarios.
+Figure 5.
+Component distribution and latency analysis across slice groups.
+Communication resources for uplink slices in Slice 1 to Slice 3 are gradually increasing.
+The middle panel statistically validates performance differences, while the rightmost panel quantifies how slice configurations shift processing time proportions among uplink, downlink, and inference operations.
+LLM services exhibit fundamentally different behavior where computational demands vary dramatically and interact complexly with network resources.
+Unlike traditional DNNs with fixed architectures processing fixed-size inputs, LLMs generate variable-length outputs through autoregressive decoding.
+The inference time depends on multiple dynamic factors, such as the input and output token count, sampling temperature, beam search width, accumulated context length, and more.
+Our measurements reveal that identical prompts can result in inference times varying by over 300% based on these generation parameters.
+More critically, this computational variability interacts with network slicing to create dynamic bottleneck migration.
+As shown in Figure 5, when network slice resources increase from minimal to abundant allocation, the system bottleneck shifts dramatically.
+Under constrained slice configurations, inference time dominates at 59.6% of total latency, as limited network resources throttle the overall pipeline.
+However, with enhanced slice resource allocation, the bottleneck migrates to transmission, which grows to 54.7% of latency.
+This migration occurs because improved network resources enable faster data transfer, exposing computational limitations that were previously masked by network constraints.
+This coupling between computation and communication invalidates traditional slice optimization approaches that assume fixed service characteristics.
+LLM services require dynamic adaptation mechanisms that consider both computational state and network conditions, adjusting resource allocation based on real-time bottleneck identification rather than static service profiles.
+2.3. Insight 3: Token Stream Unpredictability
+Traditional DNN applications generate highly predictable network traffic.
+Classification models produce fixed-size outputs (Krizhevsky et al., 2017); object detectors generate bounded detection lists (Zhao et al., 2019; Redmon et al., 2016); even video streams maintain consistent bitrates through rate control (Om et al., 2022; Alsmirat et al., 2023).
+This predictability enables efficient resource allocation through statistical multiplexing and quality-of-service guarantees based on well-characterized traffic models.
+LLM token streams violate these assumptions through three unique characteristics (Stojkovic et al., 2025).
+First, the generative nature creates inherent unpredictability in output length.
+A simple query may yield trigger a terse response or an elaborate explanation, depending on subtle prompt variations or sampling randomness.
+Second, token generation exhibits extreme burstiness as the model alternates between rapid token transmission and computational pauses for attention calculation.
+Third, multi-turn interactions introduce state dependencies where previous exchanges influence subsequent resource requirements through context accumulation.
+Figure 6.
+Temporal analysis of PRB allocation patterns across three implementation scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom).
+The x-axis represents timestamps in microseconds, while the y-axis denotes current resource blocks allocated to the UE.
+The visualization demonstrates the effectiveness of the slicing architecture in providing differentiated resource allocation.
+Figure 7.
+Temporal analysis of actual byte transmission across three scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom).
+Despite controlled PRB allocation, actual byte transmission shows high variance, particularly pronounced in LLM traffic patterns.
+Our measurements, visualized in Figures 6 and 7, reveal the practical implications of this unpredictability.
+Despite controlled Physical Resource Block (PRB) allocation following defined slice policies, actual byte transmission shows high variance and non-linear relationships with allocated resources.
+This non-linearity is particularly pronounced in LLM traffic compared to conventional data flows, as the stochastic generation process creates transmission bursts that poorly align with periodic resource scheduling.
+The unpredictability extends beyond simple variance.
+Token streams exhibit complex temporal correlations as the model generates semantically coherent sequences, creating traffic patterns that defy traditional Markovian models.
+Resource allocation strategies based on average behavior or peak provisioning both prove inadequate: average-based allocation suffers from frequent deadline violations during generation bursts, while peak provisioning wastes resources during computation-bound phases.
+Effective LLM service delivery requires fine-grained monitoring and adaptive resource management that responds to real-time generation patterns rather than statistical traffic models.
+These insights reveal why LLM services demand new wireless designs: bidirectional asymmetry requires direction-aware scheduling; computation-communication coupling needs dynamic slice optimization; and token stream unpredictability calls for adaptive resource management.
+Unlike traditional DNN applications, these characteristics necessitate rethinking wireless architectures.
+The following section presents WiLLM, our platform addressing these challenges through targeted design principles and implementation choices.
+3. WiLLM System Architecture
+The insights from Section 2 drive WiLLM’s architectural decisions.
+This section presents our design philosophy, system overview, and key innovations that collectively enable efficient LLM service delivery over wireless networks while democratizing access for researchers without wireless expertise.
+3.1. Design Philosophy
+WiLLM’s design philosophy centers on making LLM services accessible to researchers while addressing their unique technical challenges.
+Based on our empirical insights and the limitations of existing systems, we establish three core principles that guide our architectural decisions.
+Principle 1: Universal Accessibility.
+Unlike existing testbeds that require specialized devices with native slicing support (Saboorian and Xiang, 2017), WiLLM must enable any standard device to leverage network slicing capabilities for fine-grained resource allocation.
+This principle drives our universal UE compatibility design through application-layer tunneling, allowing commodity devices to access slice-specific resources that guarantee LLM service requirements.
+By bridging the gap between slice-unaware devices and slice-enabled networks, we empower standard smartphones and laptops to benefit from differentiated QoS provisioning, ensuring predictable latency and bandwidth allocation crucial for interactive LLM applications.
+This democratizes advanced network resource management beyond specialized research equipment.
+Principle 2: Dynamic Adaptability.
+The complex interplay between LLM computational variability and network conditions demands systems that can adapt in real-time.
+Static configurations fail to address the migrating bottlenecks and changing resource requirements inherent in LLM services.
+This principle motivates our dynamic slice compatibility mechanisms, multi-UE multi-slice scheduling capabilities, and dual-mode resource allocation, enabling the system to respond to varying workload characteristics without manual reconfiguration.
+Principle 3: Cross-Layer Observability.
+The tight coupling between application-layer token generation, network-layer transmission, and computation-layer inference requires visibility and control across all system layers.
+Traditional layered architectures that hide internal state prevent the coordinated optimization necessary for LLM services.
+Real-time adaptation to LLM’s dynamic behavior requires understanding not just what happens at each layer, but how events cascade across layers to impact end-to-end performance.
+This principle drives our comprehensive cross-layer API framework and monitoring infrastructure, exposing system state while maintaining appropriate abstractions for non-experts.
+These principles balance technical requirements from our measurements with practical research needs.
+Rather than prescribing specific algorithms, WiLLM provides an open platform for deploying and evaluating arbitrary scheduling strategies, removing barriers that have limited innovation at the intersection of wireless networks and LLM services.
+3.2. Key System Innovations
+3.2.1. Core Network GPU Deployment
 Figure 8.
@@
 Right: Proposed core network LLM service deployment architecture with GPU resources at network convergence points.
-LLMs as transformative technologies in wireless communications call for fundamental network architecture reconceptualization.
-Despite edge computing’s partial success in LLM deployment, its device-centric approach inherently limits scalability and optimization.
-We propose a novel core network paradigm that positions LLM resources at the nexus of high-capacity backhaul and wireless access networks, enabling comprehensive computational-communication resource integration that transcends traditional edge computing constraints.
-3.1. Deployment Paradigm for LLM Services
-Current LLM service deployment architectures exhibit fundamental architectural limitations that constrain their performance envelopes.
-As illustrated in Figure 2, traditional cloud-based configurations necessitate traversal of multiple network boundaries: from UE through gNB and CN to remote LLM servers via public internet infrastructure.
-This architectural paradigm introduces non-deterministic network latency due to Internet Gateway traversal, unpredictable congestion during peak utilization periods, and critically, a decoupling of computational and communication resource management planes that prevents cross-layer optimization.
-[…]
-The technical implementation comprises three principal architectural components:
-  1. (1)
-    Deterministic Service Path Integration: As shown in Figure 2, our implementation establishes a direct service flow from UE through gNB to the CN-GPU integrated infrastructure.
-This configuration eliminates the indeterministic routing and potential congestion points inherent in public internet traversal, enabling guaranteed service-level agreement enforcement through contiguous administrative control of the entire transmission path.
-  2. (2)
-    Hierarchical Slice-Based Resource Orchestration: The implementation leverages the slice management mechanisms detailed in Section 3.3 to establish fine-grained coordination between communication and computational resources.
-Rather than treating these as separate resource domains (as in conventional implementations), our architecture implements a unified resource allocation framework where slice policies govern both radio resource allocation at the gNB and computational resource distribution at the CN-GPU nexus.
-  3. (3)
-    End-to-End Service Flow Management: The architecture implements a comprehensive flow control mechanism that extends from UE request initiation through transmission scheduling, computational resource allocation, and response delivery.
-This coordinated approach enables advanced optimization techniques such as predictive resource allocation, where the system anticipates computational requirements based on request characteristics and proactively configures both network and computational resources accordingly.
-3.2. Transcending Edge Computing Limitations
-The core network deployment paradigm for LLM services introduces fundamental architectural, operational, and infrastructural enhancements that transcend edge computing’s intrinsic limitations.
-The hierarchical “Tree-Branch-Fruit” slicing framework (Section 3.3) enables differentiated, SLA-governed LLM orchestration, surpassing edge computing’s localized resource allocation.
-The cross-layer optimization framework (Sections 4.2) synergistically coordinates resources across network strata, realizing dynamic adaptation unattainable through edge computing’s device-centricity.
-Strategic LLM resource placement at the backhaul-RAN confluence forges a computational-communication symbiosis, transcending edge deployments’ latency-processing tradeoffs.
-Empirical insights from the WiLLM dataset (Sections 5) corroborate the core network paradigm’s transformative potential, illuminating performance benefits and optimization opportunities eclipsing edge computing.
-This paradigm establishes the foundation for the “Tree-Branch-Fruit” slicing architecture, facilitating a technical environment where both communication resources and computational capabilities can be allocated through unified slice policies.
-The implementation enables an evolutionary deployment path that preserves investment in existing infrastructure while advancing sophisticated LLM service delivery models.
-3.3. “Tree-Branch-Fruit” Slicing Architecture
-Figure 3. Tree-Branch-Fruit network slicing architecture with hierarchical service organization.
-Figure 4. Slice function architecture of WiLLM with integrated UE, gNB, Core Network, and Edge Server Components (Liu et al., 2024).
-Building upon our core network deployment paradigm, we propose a novel “Tree-Branch-Fruit” slicing architecture that transforms conventional network slicing approaches into a comprehensive service provisioning platform.
-This multi-tier architecture, illustrated in Figure.
-3, establishes a hierarchical framework that maintains backward compatibility with existing slicing mechanisms while introducing specialized LLM service slices.
-In this architecture:
-  • Tree (gNB): Represents the foundational radio access network infrastructure, serving as the common entry point for all network traffic and the root of the slicing hierarchy.
-  • Branch Slices: Correspond to traditional communication service slices defined in 5G standards, including enhanced Mobile Broadband (eMBB), Ultra-Reliable Low-Latency Communication (URLLC), and massive Machine Type Communication (mMTC).
-These branch slices are not the focus and implementations of this work.
-These branches maintain backward compatibility with existing network functions.
-  • Fruit Slices: Denote specialized LLM service slices that extend from branch slices, each configured for specific application domains and computational requirements.
-Figure.
-3 presents illustrative examples of potential fruit slice configurations, including various LLM service use cases with different model sizes tailored to specific application requirements.
-The examples illustrated in Figure.
-4 demonstrate how different application domains might be served by appropriately sized language models, with computationally intensive applications connected to branches supporting higher bandwidth and more powerful models, while lightweight IoT applications utilize smaller, more efficient models.
-These examples are not prescriptive implementations but rather illustrative use cases demonstrating the flexibility of the proposed architecture.
-This architecture establishes a “slice-as-a-LLM-service” paradigm that facilitates efficient multi-user and multi-slice coordinated scheduling through cross-layer resource optimization.
-The hierarchical structure enables:
-  • Differentiated Service Provisioning: Telecom operators can provide LLM services with varying performance guarantees and computational capacities tailored to specific application requirements and price points.
-  • Resource Isolation: The architecture enforces strict isolation between service slices, preventing performance degradation in one service from affecting others and enabling guaranteed quality-of-service levels.
-  • Domain-Specific Optimization: Each fruit slice can be independently optimized for its specific application domain, with tailored model selection, inference parameters, and resource allocation strategies.
-  • Modular Service Evolution: New LLM services can be deployed as additional fruit slices without disrupting existing services, facilitating incremental innovation.
-Our architecture achieves forward compatibility via application-layer tunneling that transcends conventional slicing constraints.
-This method encapsulates LLM traffic within standard application protocols, enabling infrastructure compatibility while maintaining architectural autonomy.
-The approach provides telecommunications operators with implementation flexibility independent of underlying radio network slice capabilities.
-This strategic decoupling of service implementation from network infrastructure allows fruit slices to evolve asynchronously from hardware upgrade cycles or standardization timelines.
-Such architectural independence ensures operational resilience across heterogeneous networks, enabling deployment of novel LLM capabilities unfettered by the limitations of conventional slicing frameworks.
-The monetary value icons associated with different fruit slices in Figure.
-3 illustrate the differentiated pricing potential of this architecture, with higher-resource slices commanding premium pricing while lightweight services remain cost-effective for mass deployment.
-This economic framework creates sustainable incentives for network operators to invest in LLM-capable infrastructure.
-The “Tree-Branch-Fruit” architecture represents an advancement over conventional slicing approaches by introducing application-specific service customization while maintaining backward compatibility with existing network functions.
-This novel paradigm establishes the foundation for our comprehensive WiLLM implementation, which will be detailed in subsequent sections.
-4. The Proposed System
-Figure 5. Comprehensive system architecture of WiLLM with UE, gNB, CN, and Edge Server Components.
-This section presents the comprehensive design and implementation of WiLLM, our novel wireless communication system specifically optimized for LLM service delivery.
-Building upon the “Tree-Branch-Fruit” slicing architecture described in Section 3, we have constructed a complete system implementation that extends the OpenAirInterface framework with specialized components for LLM service provisioning.
-Figure.
-5 illustrates the system architecture, encompassing UE, gNB, core network, and edge server components with their associated interconnections and functional modules.
-4.1. System Architecture Overview
-WiLLM adopts an innovative vertical integration architecture design, achieving end-to-end coordination from user applications to core networks through a hierarchical management framework.
-This architecture, centered on layered slice management, organically integrates four key subsystems: the CN subsystem built on the Open5GS framework, with enhanced edge computing capabilities through a dedicated Edge Server Bridge; the gNB subsystem that implements enhanced “Tree-Branch-Fruit” slice management mechanisms based on OpenAirInterface; the UE subsystem that integrates configuration management, slice control, and performance measurement function modules; and the edge server subsystem that carries various LLM deployments and executes corresponding slice policies.
-These subsystems are not simply stacked together but are tightly coupled through a hierarchical slice management framework, enabling precise scheduling and collaborative optimization of cross-layer resources, thereby ensuring the efficiency of the LLM communication system.
-4.2. Implementation Framework
-WiLLM extends OpenAirInterface with strategic innovations that balance implementation efficiency with transformative functionality.
-This approach leverages established communication foundations while introducing specialized components that fundamentally transform system capabilities:
+WiLLM fundamentally reimagines LLM service deployment by positioning GPU resources within the core network rather than at edge nodes or remote clouds.
+As illustrated in Figure 8, this architectural choice addresses the unique challenges of LLM services while leveraging existing telecommunications infrastructure.
+Traditional cloud deployments (left side of Figure 8) require traffic to traverse the public internet, introducing variable latency, unpredictable congestion, and loss of operator control over service quality (Chen et al., 2024).
+Edge deployments attempt to reduce latency but face severe limitations: insufficient computational capacity for large language models, thermal constraints in compact edge nodes, and coordination challenges across distributed edge sites (He et al., 2024; Cai et al., 2024; Nelson and Toloui, 2023; Tang et al., 2023).
+Our core network deployment (right side of Figure 8) positions GPU clusters at the convergence point between high-capacity backhaul networks and the radio access network.
+This location offers unique advantages for LLM services.
+First, it provides access to substantial computational resources through dedicated GPU infrastructure connected via high-speed interconnects.
+Second, it maintains low latency by avoiding internet traversal while serving multiple base stations from a centralized location.
+Third, it enables coordinated resource management as both communication and computation resources fall under unified administrative control.
+This deployment paradigm particularly benefits from the bidirectional asymmetry of LLM services.
+During uplink-heavy inference, the proximity to radio networks minimizes data transfer delays while powerful GPUs handle computational bottlenecks.
+For downlink-heavy generation, the high-bandwidth backhaul connections efficiently distribute generated content.
+The centralized nature also facilitates resource sharing across multiple concurrent LLM sessions, improving utilization compared to distributed edge deployments.
+3.2.2. Tree-Branch-Fruit Slicing Architecture
+Figure 9.
+Tree-Branch-Fruit network slicing architecture with hierarchical service organization.
+The tree represents base infrastructure, branches correspond to traditional service categories, and fruits denote specialized LLM services with differentiated pricing models.
+Building upon core network GPU deployment, WiLLM introduces the Tree-Branch-Fruit slicing architecture that extends traditional network slicing for LLM service differentiation.
+As depicted in Figure 9, this three-tier hierarchy enables flexible service provisioning while maintaining compatibility with existing 5G slicing mechanisms.
+The Tree layer represents the foundational radio access network infrastructure shared across all services.
+This includes the physical base stations (gNBs), spectrum resources, and basic connectivity functions.
+The tree provides common entry points for all traffic while implementing isolation mechanisms to prevent interference between different service branches.
+The Branch layer corresponds to traditional 5G network slices: enhanced Mobile Broadband (eMBB), Ultra-Reliable Low-Latency Communication (URLLC), and massive Machine Type Communication (mMTC).
+These branches maintain standard 3GPP-defined behaviors, ensuring backward compatibility with existing devices and services.
+Each branch implements specific quality-of-service parameters, resource allocation policies, and isolation guarantees as specified in 5G standards.
+The Fruit layer introduces LLM-specific service slices that extend from appropriate branches based on their primary characteristics.
+Each fruit slice is configured for particular LLM applications with tailored resource allocations, model selections, and pricing tiers.
+For instance, a real-time translation service might extend from the URLLC branch with a lightweight model ensuring low latency, while a creative writing assistant could branch from eMBB with a larger model prioritizing quality over speed.
+This hierarchical architecture enables several key capabilities.
+Service differentiation allows operators to offer multiple LLM service tiers with varying performance guarantees and pricing models, as indicated by the monetary symbols in Figure 9.
+Dynamic fruit creation permits rapid deployment of new LLM services without modifying core network functions.
+Resource isolation ensures that computational or communication demands from one LLM service do not impact others.
+Most importantly, the architecture maintains a clear evolutionary path where existing networks can progressively add LLM capabilities through fruit slice deployment.
+Figure 10.
+Comprehensive system architecture of WiLLM with UE, gNB, CN, and LLM Server Components.
+The diagram illustrates data flows, control pathways, and the integration of LLM-specific components within the telecommunications infrastructure.
+WiLLM implements five key innovations that collectively address the unique challenges of LLM service delivery while maintaining accessibility for researchers, as illustrated in the comprehensive system architecture shown in Figure 10.
 3.2.3. Dynamic Slice Compatibility
-As illustrated in Figure.
-5, our implementation extends conventional network slicing capabilities with LLM-specific service slices while maintaining backward compatibility.
-The Slice Manager of gNB encompasses both Branch Slice-UE Mapping and Fruit Slice-UE Mapping components, with bidirectional data flows (indicated by the UE Update and UE State Report pathways) enabling dynamic reassignment of resources in response to changing service demands.
-The implementation employs a virtualization layer that abstracts underlying slice implementation details, while the Branch Slice Policy Executor and Fruit Slice Policy Executor enforce service-specific policies at the gNB level.
-As shown in the left section of Figure.
-5, these policy executors interact directly with the PRB Manager, creating a decision hierarchy that translates high-level service policies into physical resource allocations.
+Traditional network slicing assumes static service characteristics (Andrade and Wickboldt, 2025), but LLM services exhibit highly variable resource requirements based on model selection, generation parameters, and user interactions.
+WiLLM’s dynamic slice compatibility extends conventional slicing with runtime adaptation capabilities.
+As shown in Figure 10, the Slice Manager components in both gNB and UE subsystems implement bidirectional communication through UE Update and UE State Report pathways.
+This enables real-time slice reconfiguration based on observed performance metrics.
+The Branch Slice Policy Executor and Fruit Slice Policy Executor translate high-level service requirements into concrete resource allocation decisions, adapting to changing computational loads and network conditions.
+The dynamic compatibility specifically addresses LLM’s variable inference times and unpredictable token generation patterns.
+When the system detects a shift from computation-bound to communication-bound operation, slice parameters automatically adjust to maintain service quality.
+This adaptation occurs transparently, requiring no manual intervention from researchers.
 3.2.4. Universal UE Compatibility
-To address the limitations of current UE implementations, which often lack native slicing capabilities, WiLLM incorporates a middleware-based slice access mechanism through the Slice Manager of UE component.
-The data flow illustrated in Figure.
-5 demonstrates how this component interfaces with both the Configuration Manager and Hot Start Module, enabling UEs without native slicing support to access specialized LLM services through application-layer tunneling techniques that encapsulate service traffic within standard data streams.
-This compatibility layer effectively bypasses the limitations of current community implementations of primary slicing, which remain incomplete and inconsistently supported across device ecosystems, thereby expanding the addressable user base for LLM services.
-4.2.3. Multi-UE Multi-Slice Scheduling Capability
-WiLLM implements a comprehensive multi-UE-multi-slice coordinated scheduling mechanism that operates through a two-phase process.
-The Scheduler module in Figure.
-5 forms the nexus of this coordination, receiving inputs from multiple sources including the Buffer Manager (which provides Buffer Status information), HARQ Manager (which signals Retransmission Requirements), and the Slice Allocation of UE component.
-The multi-UE-multi-slice coordination operates through a two-phase process:
-  • A global allocation phase distributes available physical resource blocks across active slices according to their relative priorities and resource guarantees.
-  • An intra-slice scheduling phase allocates resources to individual UEs within each slice based on service-specific metrics.
-This hierarchical approach ensures fair and efficient resource utilization while maintaining service isolation, effectively enabling legacy devices to benefit from slice-specific performance guarantees without hardware modifications.
+A critical barrier in existing platforms is the requirement for specialized UE equipment with native slicing support.
+WiLLM eliminates this barrier through application-layer tunneling mechanisms that enable any standard device to access LLM services.
+The Slice Manager of UE component in Figure 10 implements a middleware layer that intercepts application traffic and encapsulates it with appropriate slice identifiers.
+This approach bypasses the limitations of current UE implementations, which often lack complete slicing support.
+The Configuration Manager and Hot Start Module work together to establish slice contexts without requiring protocol-level modifications.
+This innovation is crucial for democratizing access to LLM service research.
+Researchers can use commercial smartphones, laptops, or IoT devices without specialized firmware or hardware modifications.
+The application-layer approach also enables rapid prototyping of new service types without waiting for standardization or device updates.
+3.2.5. Multi-UE Multi-Slice Scheduling
+LLM services create complex scheduling challenges when multiple users with different service requirements share network resources.
+WiLLM implements sophisticated scheduling mechanisms that coordinate resources across multiple UEs and slice types simultaneously.
+The Scheduler module at the center of Figure 10 receives inputs from multiple sources: Buffer Status from the Buffer Manager, Retransmission Requirements from the HARQ Manager, and Slice Allocation policies from the mapping components.
+This multi-dimensional optimization problem is solved through a hierarchical approach: first allocating resources across slices based on their aggregate requirements, then scheduling individual UEs within each slice based on service-specific metrics.
+The scheduling algorithm specifically accounts for LLM characteristics.
+Token generation bursts receive priority handling to minimize user-perceived latency, while background model updates are scheduled during resource valleys.
+The system maintains fairness across users while respecting the differentiated service levels promised by different fruit slices.
 3.2.6. Dual-Mode Resource Scheduling
-WiLLM implements complementary “separated” and “embedded” resource scheduling methodologies to accommodate diverse operational requirements and infrastructure constraints:
-  • Embedded Scheduling Mode: The embedded approach implements slice-aware resource allocation directly within the scheduler pipeline, as formalized in Algorithm 1.
-Figure.
-5 illustrates how this approach integrates with the Resource Grid and Spectrum Usage Tracking components to maintain awareness of physical layer constraints while implementing policy decisions.
-  • Separated Scheduling Mode: The separated approach decouples slice-specific policy decisions from core scheduling logic, employing an external decision engine that solves a global optimization problem and subject to resource constraints and slice isolation requirements.
-The data flow in Figure.
-5 shows how the Resource Update pathway enables this external engine to maintain synchronization with the core scheduling system.
-The interaction between the Uplink/Downlink Allocation Request pathways and the PUCCH Allocator modules for both uplink and downlink control channels (as depicted in Figure.
-5) demonstrates how the system coordinates the allocation of scarce control channel resources across multiple UEs and slices.
+To balance implementation efficiency with optimization sophistication, WiLLM implements complementary embedded and separated scheduling modes, each suited to different operational scenarios.
+In embedded mode, slice-aware resource allocation integrates directly within the main scheduling pipeline.
+This provides low-latency adaptation suitable for real-time LLM interactions.
+The scheduler modifies its decisions based on slice policies without external intervention, enabling microsecond-scale adaptations to changing conditions.
+In separated mode, an external optimization engine analyzes longer-term patterns and computes globally optimal resource allocations.
+This engine interfaces through the Resource Update pathway, enabling sophisticated algorithms including machine learning-based prediction and multi-objective optimization.
+The separated mode excels at identifying persistent patterns in LLM usage and pre-allocating resources accordingly.
+The dual-mode design allows researchers to experiment with different scheduling strategies without modifying core system components.
+Simple policies can be implemented in embedded mode for immediate testing, while complex algorithms can be developed in separated mode with full access to historical data and predictive models.
 3.2.7. Cross-Layer API Framework
-WiLLM introduces a comprehensive API architecture (detailed in Figure.
-16) that facilitates vertical integration between application requirements and network-level resource allocation.
-As shown in Figure.
-5, this framework enables multiple feedback pathways:
-  1. (1)
-    Performance Feedback Loop: The UE State Report, Configuration/Feedback/Buffer, and UE Update connections establish a closed-loop feedback system that continuously monitors service performance and adapts resource allocation accordingly.
-The bidirectional nature of these connections, clearly depicted in Figure.
-5, enables adaptive behavior at multiple timescales.
-  2. (2)
-    Service-Specific Optimization: The LLM Service ID mechanism enables adaptation of network parameters based on computational characteristics of different LLM implementations.
-Figure.
-5 illustrates how this information flows from the Edge Server Subsystem through the Slice Manager to influence allocation decisions.
-  3. (3)
-    Resource Utilization Monitoring: The Resource Usage and Slice Allocation connections provide visibility into consumption patterns, establishing a cross-layer information flow that informs both immediate scheduling decisions and longer-term policy refinements.
-4.2.6. Deployment Flexibility
-The system supports flexible LLM deployment configurations, with the Edge Server Large Language Model Deployment component hosting multiple LLM variants with different parameter sizes as shown in Figure.
-5, and the CN-Edge Server Bridge facilitating seamless integration between computational resources and network infrastructure through standardized file and network interfaces.
-4.3. The Database Module
-A distinctive feature of WiLLM is its integrated data collection module (shown in the upper section of Figure.
-5), which continuously refines resource allocation strategies based on operational metrics.
-This module creates a meta-feedback loop that transcends individual communication sessions, collecting performance data from UE’s Performance, gNB Measurement, and Server Measurement components to inform the evolution of the Customized QoS Scheduler and Model components.
-The system incorporates a comprehensive Database component (illustrated in the right section of Figure.
-5) that persistently stores operational metrics, enabling longitudinal analysis of system performance.
-This data repository serves as both an analytical resource for offline optimization and a training corpus for online learning algorithms.
-4.4. System Validation Infrastructure
-To validate practical performance, we deployed a fully functional testbed (Figure.14 in Appendix) integrating a complete wireless communication chain with LLM service deployment.
-Following the logical architecture shown in Figure.
-5, the core computing infrastructure employs an Intel Core i9-14900KF processor as gNB host, USRP B210 for RF frontend, and NVIDIA RTX 4090 GPU for LLM inference, supporting multi-user, multi-slice concurrent testing scenarios.
-The system’s operational status can be monitored through a custom GUI (Figure.
-15 in Appendix).
-This configuration achieves complete system integration from radio access to core network, providing an ideal environment for evaluating WiLLM in practical deployments.
-5. The WiLLM Dataset
-We present the first comprehensive LLM communication dataset with synchronized metrics, enabling reproducible research and systematic evaluation.
-Appendix F details components, acquisition, synchronization, and error descriptions.
-5.1. Dataset Metrics Overview
-Our dataset contains 1,649,996 records with 58 metrics spanning the LLM service delivery chain.
-It covers four scenarios: static UE and gNB (290,653 records), dynamic UE and static gNB (363,906 records), static UE and dynamic gNB (430,369 records), and dynamic UE and gNB (565,068 records).
-“Dynamic” indicates dynamically allocated network slices, while “static” indicates invariant parameters throughout measurements.
-The dataset includes 22 metrics at the user equipment layer, 25 at the wireless access network layer, and 18 at the edge server layer.
-Figures 17 and 18 confirm stable base station power and noise levels during data collection.
-We implemented NTP-based distributed synchronization, calibrating all devices through a common server for microsecond-level sampling precision.
-By recording timestamps at both the UE and server endpoints and using latency compensation algorithms, we maintained synchronization errors within $\pm$1.0 milliseconds.
-5.2. Metric Architecture
-The dataset provides synchronized metrics across three architectural layers: UE Metrics (Appendix Table 4) capture end-user experience through latency, timing, resolution, modalities, and payload characteristics; gNB Metrics (Appendix Table 6) document wireless conditions via radio parameters (MCS, CQI, BLER), resource allocation, transmission performance, and slice configuration; and CN with Edge Server Metrics (Appendix Table 5) measure computational aspects through inference timing, token processing, model loading, and response generation.
-The dataset’s value lies in its temporal synchronization, enabling precise correlation between user experience, network conditions, and computational performance.
-This facilitates advanced research: developing ML models for scheduling, implementing data-driven allocation, creating cross-layer analytics, and validating theoretical models.
-This corpus establishes a foundation for reproducible research in wireless systems supporting LLM services.
-5.3. Preliminary Findings and Conclusions
+To enable effective research without requiring wireless expertise, WiLLM provides comprehensive APIs that abstract network complexity while exposing necessary control and monitoring capabilities.
+As shown in the online Figure 16, the API architecture implements a three-tier structure.
+The User Management API handles service registration, authentication, and preference management.
+Researchers interact with familiar application-level concepts like ”model selection” and ”latency requirements” rather than radio network parameters.
+The API automatically translates these high-level specifications into appropriate network configurations.
+The System Management API provides slice lifecycle management and service orchestration.
+Through RESTful interfaces, applications can discover available services, request specific configurations, and monitor performance metrics.
+The API includes WebSocket endpoints for real-time monitoring of token generation rates, latency breakdowns, and resource utilization.
+The Resource Management API offers fine-grained control for advanced users who need to optimize specific aspects of service delivery.
+This includes interfaces for custom scheduling policies, resource reservation, and predictive allocation based on application hints.
+Importantly, this API remains optional—basic LLM services function without any low-level configuration.
+The cross-layer design ensures information flows efficiently between system components.
+Token generation patterns from the application layer inform network scheduling decisions, while network congestion indicators trigger application-layer adaptation.
+This bidirectional information flow, illustrated by the various feedback pathways in Figure 10, enables holistic optimization impossible in traditional layered architectures.
+These five innovations work synergistically to create a system that is both powerful enough to handle LLM’s unique challenges and accessible enough for researchers without wireless expertise.
+The following section details how we realized these concepts through concrete implementation on the OpenAirInterface platform.
+4. Implementation
+Transforming WiLLM’s architectural vision into a functional system required substantial engineering effort.
+This section details our implementation approach, focusing on how we extended OpenAirInterface to support LLM services while maintaining compatibility and usability.
+4.1. Building on OpenAirInterface
+We selected OpenAirInterface (OAI) as our implementation foundation for several compelling reasons.
+First, OAI provides a complete, modular implementation of 4G/5G protocol stacks from physical layer to core network, enabling end-to-end system modifications.
+Second, its open-source nature allows deep modifications necessary for LLM-specific optimizations.
+Third, OAI’s active community and extensive documentation lower the barrier for researchers to understand and extend our work.
+However, implementing WiLLM required substantial extensions beyond OAI’s baseline capabilities.
+Our modifications span three critical areas:
+Slice Management Extensions.
+While OAI includes basic network slicing support, it lacks the dynamic adaptation required for LLM services.
+We extended the Radio Resource Control (RRC) and Service Data Adaptation Protocol (SDAP) layers to support runtime slice reconfiguration.
+The modifications enable slice parameters to be adjusted based on real-time performance metrics without disrupting ongoing sessions.
+We implemented new signaling procedures that propagate slice updates from the core network through the RAN to individual UEs, maintaining consistency across the system.
+Scheduler Enhancements.
+The OAI scheduler originally implements proportional fair and round-robin algorithms optimized for traditional traffic patterns.
+We augmented it with LLM-aware scheduling logic that recognizes token generation patterns and adapts resource allocation accordingly.
+Our modifications include: (1) traffic classification modules that identify LLM flows based on packet patterns, (2) dual-mode scheduling support with hooks for external optimization engines, and (3) direction-aware resource allocation that treats uplink and downlink asymmetrically based on identified bottlenecks.
+Cross-Layer Information Flow.
+Traditional OAI maintains strict layer separation, preventing the coordinated optimization necessary for LLM services.
+We implemented information conduits that enable controlled cross-layer communication while preserving architectural modularity.
+New interfaces expose internal state from MAC, RLC, and PDCP layers to higher-level orchestration components.
+Similarly, application-layer hints about token generation patterns flow down to influence scheduling decisions.
+These modifications required careful design to avoid violating layer abstractions while enabling necessary information exchange.
+4.2. Engineering Challenges and Solutions
+Implementing WiLLM’s architectural innovations presented several significant engineering challenges that required novel solutions based on the unique characteristics of LLM services identified in Section 2.
+Challenge 1: Universal UE Compatibility Without Protocol Modifications.
+A critical barrier in existing platforms is the requirement for specialized UE equipment with native slicing support.
+Current community implementations of network slicing remain incomplete and inconsistently supported across device ecosystems, effectively excluding standard commercial devices from accessing advanced network features.
+Why existing systems fail: Current platforms like OAI and srsRAN hard-code slice support at the protocol level, requiring modifications to UE protocol stacks for slice access.
+This approach excludes the vast majority of commercial devices and creates an impractical barrier for researchers who lack the capability to modify device firmware or develop custom UE implementations.
+Solution: We developed a middleware-based slice access mechanism through application-layer tunneling techniques.
+The Slice Manager of UE component implements a compatibility layer that encapsulates LLM service traffic within standard application protocols, enabling infrastructure compatibility while maintaining architectural autonomy.
+This approach provides telecommunications operators with implementation flexibility independent of underlying radio network slice capabilities, allowing fruit slices to evolve asynchronously from hardware upgrade cycles or standardization timelines.
+Challenge 2: Dynamic Slice Reconfiguration for Variable LLM Workloads.
+LLM services exhibit highly variable resource requirements based on model selection, generation parameters, and user interactions, as demonstrated in our empirical analysis.
+Traditional network slicing assumes static service characteristics, creating a fundamental mismatch with LLM’s dynamic nature.
+Why existing systems fail: Traditional slicing implementations in platforms like FlexRIC and Open5GS follow 3GPP specifications that assume static service profiles.
+Once configured, slice parameters remain fixed throughout the session lifetime.
+This rigid approach cannot accommodate LLM’s dynamic resource requirements that vary by orders of magnitude within a single session.
+Solution: We extended conventional slicing with runtime adaptation capabilities through bidirectional communication pathways (UE Update and UE State Report) that enable real-time slice reconfiguration based on observed performance metrics.
+The Branch Slice Policy Executor and Fruit Slice Policy Executor translate high-level service requirements into concrete resource allocation decisions, adapting to changing computational loads and network conditions without disrupting ongoing sessions.
+Challenge 3: Cross-Layer Coordination for Computation - Communication Coupling.
+Our measurements revealed tight coupling between LLM computational variability and network performance, where slice configuration changes cause dynamic bottleneck migration between computation and communication.
+Traditional layered architectures prevent the coordinated optimization necessary for managing this coupling.
+Solution: We implemented a comprehensive cross-layer API framework that facilitates vertical integration between application requirements and network-level resource allocation.
+The system includes performance feedback loops, service-specific optimization pathways that enable LLM Service ID information to influence network decisions, and resource utilization monitoring that provides visibility into consumption patterns across system layers while maintaining appropriate abstractions for non-experts.
+Challenge 4: Multi-UE Multi-Slice Scheduling Complexity.
+LLM services create complex scheduling challenges when multiple users with different service requirements share network resources.
+The variable inference times and unpredictable token generation patterns identified in our analysis significantly complicate resource allocation compared to traditional traffic.
+Solution: We implemented a hierarchical scheduling mechanism that operates through a two-phase process: a global allocation phase distributes available physical resource blocks across active slices according to their relative priorities and resource guarantees, followed by an intra-slice scheduling phase that allocates resources to individual UEs within each slice based on service-specific metrics.
+This approach ensures fair and efficient resource utilization while maintaining service isolation.
+Challenge 5: Temporal Synchronization Across Distributed Components.
+Correlating LLM service events across UE, RAN, and CN components requires precise time synchronization, particularly challenging given the variable processing delays inherent in LLM inference and the need for microsecond-level accuracy in performance analysis.
+Solution: We implemented NTP-based distributed synchronization across all system components, calibrating devices through a common server to achieve microsecond-level sampling precision.
+By recording timestamps at both UE and server endpoints and implementing latency compensation algorithms, we maintained synchronization errors within ±1.0 milliseconds.
+This precision enables the accurate cross-layer correlation analysis that forms the foundation of our empirical insights.
+These engineering solutions demonstrate that WiLLM’s architectural innovations can be realized through practical implementations that address the unique challenges of LLM service delivery while maintaining system stability and research accessibility.
+5. WiLLM in Action: Deployment, Dataset, and Case Study
+This section demonstrates WiLLM’s practical capabilities through comprehensive deployment, data collection, and real-world application development.
+Unlike conventional wireless testbeds that focus on protocol compliance, WiLLM enables researchers to study and optimize LLM services across realistic wireless environments.
+5.1. Real-World Deployment
+WiLLM’s operational deployment demonstrates the system’s capability to deliver end-to-end LLM services while providing a robust experimentation platform for researchers.
+System Performance and Capabilities: As illustrated in our online appendix Figure 14, our deployed testbed achieves full end-to-end LLM service delivery with commercial devices.
+The system successfully handles concurrent multi-user sessions with differentiated service levels through the Tree-Branch-Fruit slicing architecture.
+The deployment maintains stable operation across diverse LLM workloads, from lightweight text generation to compute-intensive multimodal processing, validating the practical viability of core network GPU deployment for production environments.
+Research Platform Features: The testbed provides researchers with unprecedented experimental control through its integrated monitoring and adaptation capabilities.
+Real-time performance visualization enables immediate observation of cross-layer interactions, while programmable APIs allow dynamic reconfiguration without service interruption.
+Researchers can experiment with various LLM models (LLaVA, Llama 3.2) and network configurations through simple parameter adjustments, without requiring telecommunications expertise.
+The platform’s stability and ease of use have enabled extended experimental campaigns lasting multiple days without manual intervention.
+Accessibility and Adoption: A key achievement is enabling AI researchers to experiment with communication resource allocation without deep protocol stack expertise.
+Through intuitive Python APIs and web interfaces, researchers can specify scheduling policies, adjust slice parameters, and control bandwidth allocation—making wireless resource management accessible to those focused on LLM optimization.
+The platform abstracts complex protocol details while exposing meaningful control knobs: AI researchers can explore how different QoS configurations, priority schemes, and resource partitioning strategies impact their models’ performance.
+Standard devices connect seamlessly through our application-layer approach, allowing experiments on commodity hardware.
+This design bridges the gap between AI and networking communities, enabling AI researchers to actively engage with communication-layer optimizations rather than treating the network as a black box.
+5.2. Dataset
+The WiLLM dataset represents the first comprehensive empirical foundation specifically designed for wireless LLM research, providing synchronized multi-layer measurements essential for algorithm development and validation.
+Dataset Architecture and Coverage: Our dataset comprises 1,649,996 synchronized records collected across 58 metrics spanning three system layers.
+The measurements cover four distinct operational scenarios with varying dynamics: static UE with static gNB (290,653 records), dynamic UE with static gNB (363,906 records), static UE with dynamic gNB (430,369 records), and dynamic UE with dynamic gNB (565,068 records).
+This systematic variation enables researchers to analyze system behavior under different mobility and resource management conditions.
+Multi-Layer Metric Synchronization: The dataset captures 22 UE-layer metrics including request latency, response timing, image resolution parameters (320×240 to 640×480 pixels), response word count (50-200 words), and payload characteristics.
+At the gNB layer, 25 metrics document radio conditions (MCS, CQI, BLER, SNR), PRB allocation patterns, slice configurations, and scheduling decisions.
+The Core/Edge layer contributes 18 metrics measuring LLM inference timing, token generation rates, model selection (LLaVA, Llama 3.2), and GPU utilization patterns.
+Through NTP-based synchronization maintaining ±1.0ms precision, these metrics enable accurate correlation between user experience and system-level events.
+Dataset Uniqueness: Unlike traditional wireless datasets focused on channel measurements or throughput statistics, WiLLM’s dataset explicitly captures the interplay between token generation patterns, wireless resource allocation, and user-perceived performance.
+The synchronized multi-layer nature reveals complex interactions impossible to observe through isolated measurements.
+The temporal synchronization enables researchers to correlate LLM inference events with network resource allocation decisions, exposing the dynamic bottleneck migration between computation and communication that characterizes LLM services.
+Reproducibility and Standardization: The dataset includes comprehensive documentation with detailed metric definitions (the online Table 4) and collection parameters, enabling proper utilization and result reproduction.
+By providing standardized measurement data, researchers can evaluate their algorithms against the same realistic conditions, ensuring fair comparison across different approaches.
+This standardization is crucial for advancing the field beyond isolated demonstrations toward systematic performance improvements.
+5.3. Case Study: Smart Glasses Application
+To demonstrate WiLLM’s practical utility for application development, we implemented a smart glasses system that validates our platform’s ability to support real-world LLM applications while addressing human-computer interaction requirements.
+5.3.1. Application Architecture and Integration
 Figure 11.
-Component distribution and latency analysis for uplink transmissions across resolution groups.
-The x-axis labels R1 - R6 represent various resolution groups for image requests sent to LLM, where R1 - R6 correspond to increasing resolutions, with “n” representing the sample size per group, and the lower value of “n” indicating the resolution range for each group.
-The left panel shows the temporal distribution of communication components (uplink, downlink, inference), the center panel presents statistical trends with standard errors, and the right panel quantifies each component’s percentage contribution to total latency, illustrating how higher resolutions affect time allocation in the transmission pipeline.
-Figure 7.
-Component distribution and latency analysis for downlink transmissions across resolution groups.
-This figure illustrates the temporal characteristics of downlink transmissions stratified by resolution groupings.
-Similar to Figure.
-6, it provides three complementary views: component distribution, statistical latency metrics, and proportional component analysis.
-All legend elements in the figure remain consistent with Figure.
-6, with the only distinction being the utilization of downlink samples.
-This section presents preliminary findings derived from our dataset, with particular focus on communication directionality characteristics and the impact of network slicing on performance metrics.
-5.3.1. Uplink vs. Downlink Characteristic Analysis
-Finding 1: Our dataset reveals distinct performance profiles for uplink and downlink communication scenarios.
-In the uplink case (UE image request $\rightarrow$ LLM inference $\rightarrow$ text response), as visualized in Figure.
-6, we observe that inference time constitutes a substantial proportion of the total processing latency (74-87% across resolution groups).
-The relative contribution of uplink transmission time marginally increases (11-25%) with the rise in resolution.
-Finding 2: In contrast, the downlink scenarios (text request $\rightarrow$ LLM text response $\rightarrow$ image response), depicted in Figure.
-7, demonstrate markedly different characteristics.
-Downlink transmission time dominates the overall latency budget (81-86%), with inference time contributing a significantly smaller proportion (12-17%).
-Conclusions: These directional performance characteristics validate WiLLM’s “Tree-Branch-Fruit” slicing architecture and LLM deployment paradigm in several critical ways:
-  1. (1)
-    In the uplink scenario, the largest delay comes from inference time, followed by uplink transmission time.
-This finding supports our strategy of deploying LLMs with GPUs in the CN.
-This is achieved by reducing the inference time through distributed LLM processing at the CN and minimizing uplink transmission delays by staying close to users, thereby avoiding the need to go through the public Internet.
-  2. (2)
-    The pronounced asymmetry between uplink and downlink transmission characteristics empirically validates our “Fruit Slice” architectural paradigm, confirming that effective monetization of LLM services requires direction-specific slice configurations that independently optimize bandwidth and computation intensive workloads, a capability that is not achievable through conventional slicing approaches.
-  3. (3)
-    The pronounced resource requirement asymmetry between uplink and downlink scenarios provides compelling empirical justification for direction-specific scheduling mechanisms.
-These findings validate the architectural necessity of dual-mode scheduling capabilities, confirming that efficient LLM service delivery requires dynamic, direction-aware resource provisioning rather than single-direction allocation policies.
-This increased flexibility in provisioning is a key enabler of the slice-based monetization framework.
-These insights reveal that WiLLM’s paradigm and architecture, substantiated by real-world datasets, demonstrate critical advantages over conventional systems in operational dimensions.
-These advancements provide a robust framework for direction-specific resource optimization that traditional wireless systems cannot achieve.
-We provide further analysis of the reasons for the experimental phenomenon and additional notes in the appendix F.1.
-5.3.2. Network Slicing Impact Analysis
-Figure 8.
-Component distribution and latency analysis across slice groups.
-In the figure, the communication resources for uplink slices in Slice 1 to Slice 3 are gradually increasing.
-This examines performance impacts of different slice configurations on communication component distribution, highlighting distinct latency profiles for three configurations.
-The middle panel statistically validates performance differences, while the rightmost panel quantifies how slice configurations shift processing time proportions among uplink, downlink, and inference operations.
-Figure 9.
-This figure presents a temporal analysis of PRB allocation patterns across three implementation scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom).
-The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_rbs”, indicating the current number of resource blocks allocated to the UE.
-The visualization demonstrates the effectiveness of the slicing architecture in providing differentiated resource allocation while maintaining threshold compliance, with slice-distinguished traffic showing clear separation between service classes.
-The area below the dashed line is labeled as “basic data” with no data transmission, while the area above corresponds to “core data”.
-Figure 10.
-The figure illustrates a temporal analysis of PRB allocation patterns across three scenarios: normal traffic (top), slice-enabled traffic (middle), and slice-distinguished traffic (bottom).
-The x-axis represents timestamps in microseconds, while the y-axis denotes “ulsch_current_bytes”, indicating the current transmitted bytes of the UE.
-Finding 3: The effectiveness of our “Tree-Branch-Fruit” slicing architecture is demonstrated through comparative analysis of network performance under different slicing configurations.
-Figure.
-8 illustrates how slice configuration affects the temporal distribution of processing components, with notable variations in inference time allocation (43.1% to 59.6%) and uplink transmission time (28.9% to 54.7%) across different slicing arrangements.
-This demonstrates that slicing indeed affects the latency of LLM communication services and its components, which supports WiLLM’s paradigm.
-Finding 4: Examining PRB allocation patterns in Figure.
-9 alongside actual byte transmission data in Figure.
-10 reveals critical insights into slicing efficiency.
-Our implementation successfully achieves differentiated PRB allocation conforming to slice policies.
-Moreover, there exists a discrepancy between theoretical resource allocation and actual transmission patterns.
-Specifically, byte traffic does not scale proportionally with PRB allocation, suggesting potential resource utilization inefficiencies within the slicing framework.
-This indicates that due to the complex interplay of elements affecting communication, system execution mechanisms, and varying UE requests, PRB and actual communication bytes are not strictly linearly correlated.
-Conclusions: These empirical observations demonstrate the critical value of WiLLM in several dimensions:
-  1. (1)
-    The significant variability in component allocation across slice configurations validates that WiLLM’s “Tree-Branch-Fruit” architecture enables effective LLM-specific service slicing, supporting our proposed diverse pricing patterns and underscoring the necessity of finer-grained resource allocation mechanisms.
-  2. (2)
-    The observed allocation-utilization gap reveals that theoretical and simulation-based approaches often fail to capture real-world complexities, confirming WiLLM’s role as an essential empirical platform bridging theory and implementation.
-  3. (3)
-    The non-linear relationship between resource allocation and actual transmission performance underscores the necessity of WiLLM’s cross-layer API framework and dual-mode scheduling capabilities, which enable adaptive optimization based on observed rather than simulated performance patterns.
-The analysis of the causes of the observed phenomenon is presented in the Appendix F.2.
-These findings substantiate WiLLM’s fundamental contribution: while traditional approaches rely on simplified resource allocation models, our system provides the empirical foundation and architectural flexibility required for truly efficient LLM service delivery in wireless environments.
-This empirical grounding enables telecommunications operators to achieve differentiated service quality and optimal resource utilization, a critical requirement for commercially viable LLM wireless services.
-5.4. Benchmark
-Analyzing datasets reveals intricate dependencies among communication resources, computational capacity, and service quality, requiring specialized metrics.
-As traditional network metrics fail to capture LLM service delivery’s multidimensional nature, we propose two complementary metrics for evaluating LLM wireless communication systems.
-The LLM-Aware Resource Efficiency Index (LAREI) quantifies the efficiency relationship between request complexity, computational demands, and resource utilization:
-$\text{LAREI}=\frac{\text{RDV}\cdot\log(1+\text{LLM\_Para})}{\text{Resources}\cdot\text{Latency}}\cdot\omega$
-Where RDV is request data volume, LLM_Para represents model complexity in billions of parameters, Resources denotes allocated communication resources, Latency indicates end-to-end response time, and $\omega$ is a normalization coefficient.
-For slicing effectiveness evaluation, we introduce the LLM Slice Efficiency Quotient (LSEQ):
-LSEQ=RDV_slice(1Error Rate)LLM_Para_sliceSlice $\text{LSEQ}=\frac{\text{RDV\_slice}\cdot(1-\text{Error Rate})\cdot\sqrt{\text{LLM\_Para\_slice}}}{\text{Slice Resources}}\cdot\delta$
-[…]
-These metrics establish a framework for evaluating and optimizing wireless LLM communication systems, balancing computational sophistication with communication efficiency.
-Detailed analysis is presented in Appendix G.
-6. A Case Study of Smart Glasses
-6.1. Smart Glasses Integrated in WiLLM
-Figure 11.
-Smart glasses hardware integrates with the WiLLM wireless LLM service, following a typical UE-gNB-CN network architecture.
-The UE consists of customizable smart glasses and a Mini PC, which also houses the 5G module.
-A camera positioned at the glabella captures images, enabling collection and querying through a custom human-computer interaction logic.
-Figure 12.
-Gesture-triggered 5G LLM smart glasses data flow.
-The smart glasses application activates through embedded gesture recognition as the trigger.
-Upon detecting a five-finger extension followed by a grasping motion, the system captures the current visual feed and transmits it as a request to the LLM.
-To validate WiLLM’s effectiveness in real-world applications, we implemented a small but quintessential case study using augmented reality smart glasses as a resource-constrained terminal device requiring LLM integration.
-Figure.
-11 illustrates the experimental setup, showing the hardware components of our custom smart glasses platform interfaced with WiLLM network infrastructure.
-The glasses feature a 5G module, computing unit, and camera capable of capturing visual data for LLM interpretation.
-This system enables users to query about scenes before them, allowing the LLM to interpret the visual context through a LLaVA model deployed in the core network host, accessed via gNB.
-Key Validation: The smart glasses implementation demonstrates a concrete manifestation of WiLLM’s “Tree-Branch-Fruit” slicing architecture, where the device functions primarily as a visual I/O interface while the computational workload is handled by GPU resources positioned at the network convergence point, as proposed in Section 3.
-Figure.
-12 shows the interactive paradigm between user and LLM service.
-The system uses gesture recognition to convert physical movements into contextual prompts for LLM processing, creating a bidirectional channel where gestures initiate queries and responses are displayed through the glasses’ interface.
-6.2. Objective from an HCI Perspective
-Human-computer interaction (HCI) research shows that interaction quality depends less on minimal latency than on consistent, predictable response patterns (Nabiyouni et al., 2017; Jacobs et al., 1997).
-Users adapt to predictable systems, but unpredictable delays disrupt interaction continuity.
-Our goal is to identify configurations that stabilize latency around 2 seconds for the smart glasses, balancing computational efficiency and perceptual continuity within wireless LLM constraints.
-Empirical Insight: This case study revealed cross-layer dependencies where UE requests influenced communication resource utilization and allocation at the gNB, which affected LLM processing at the core network.
-This substantiates the complex relationships between resource allocation and performance documented in Section 5.
-6.3. Approaches to Slice Optimization
-Figure 13. Online learning progress of the slice selection task for the smart glasses. Gradually converge to the target slice.
-We deployed two complementary methodologies to determine optimal slice configuration.
-Our offline approach utilized WiLLM’s data collection capabilities to analyze six slice configurations, identifying those that consistently maintained the 2-second latency target with minimal variance.
-Concurrently, we implemented an adaptive online learning framework using the UCB algorithm that continuously monitored performance metrics and dynamically adjusted slice selection.
-As shown in Figure.
-13, this method demonstrated progressive refinement toward optimal configuration with reduced latency fluctuations after convergence.
-While offline methods provided thorough validation requiring extensive initial data, online learning enabled adaptive optimization with initial exploration costs.
-Both approaches effectively identified suitable configurations, demonstrating WiLLM’s cross-layer optimization capabilities for resource-constrained applications.
-Empirical Insight: The smart glasses case study provides empirical evidence that LLM service quality in wireless environments, particularly in some wearable HCI applications, depends more on maintaining predictable latency through intelligent slicing rather simply maximizing resources.
-While this single case offers only a directional indicator for resource-constrained applications, it demonstrates that WiLLM’s architectural principles effectively support interactive LLM experiences on computationally limited devices.
-Our dual optimization methodologies illustrate WiLLM’s availability of the dataset and cross-layer optimization potential, enabling sophisticated resource allocation tailored to specific application requirements.
-The hardware configurations and software components have been open-sourced to facilitate replication and expansion into other application areas.
+Smart glasses integrated with WiLLM.
+The system consists of customized smart glasses paired with a compact processing unit housing the 5G module, demonstrating practical deployment of our Tree-Branch-Fruit architecture.
+Our smart glasses implementation, illustrated in Figure 11, represents a quintessential validation of WiLLM’s design principles.
+The hardware comprises custom-built smart glasses with an integrated camera positioned at the glabella for natural field-of-view capture, paired with a compact processing unit containing a 5G module and minimal computational resources.
+This configuration validates our architectural principle that resource-constrained devices can access sophisticated LLM capabilities through core network deployment.
+The system architecture follows WiLLM’s design precisely: glasses function as intelligent sensors and display interfaces while computational processing occurs at GPU-equipped core network infrastructure.
+This separation validates our decision to position LLM resources at network convergence points rather than attempting resource-intensive edge deployment.
+Figure 12. Gesture-triggered interaction paradigm demonstrating seamless integration between physical actions and LLM service delivery through WiLLM’s bidirectional communication capabilities.
+The interaction paradigm, shown in Figure 12, demonstrates practical implementation of our bidirectional communication architecture.
+Users trigger scene analysis through intuitive gestures—extending five fingers followed by grasping motion—which captures visual frames and initiates LLM queries.
+This workflow validates WiLLM’s capability to handle the asymmetric traffic patterns identified in Section 2: high-resolution images flow uplink for scene understanding while generated descriptions return downlink for display.
+5.3.2. Human-Computer Interaction Optimization
+Performance Objectives: Human-computer interaction research establishes that user satisfaction depends more on predictable response patterns than absolute minimal latency  (Jacobs et al., 1997; Nabiyouni et al., 2017).
+This principle guided our optimization target: achieving stable 2-second response latency rather than pursuing minimal but variable delays.
+This objective balances perceptual continuity, technical feasibility, and system reliability within realistic network conditions.
+Optimization Methodologies: We employed complementary approaches demonstrating WiLLM’s versatility for algorithm development:
+Offline Analysis: Leveraging WiLLM’s comprehensive data collection capabilities, we systematically evaluated six slice configurations across extended operational periods.
+Each configuration was tested with diverse visual scenes and varying network conditions to build statistical performance models.
+Analysis revealed that intermediate slice allocations provided optimal balance between resource efficiency and performance consistency.
+Online Learning: We implemented an Upper Confidence Bound (UCB) algorithm that continuously optimizes slice selection based on observed performance.
+The algorithm initially explores different configurations, gradually converging toward those consistently meeting latency targets.
+Figure 13. Online learning progress for slice selection optimization, demonstrating WiLLM’s support for adaptive algorithm development and validation.
+As shown in Figure 13, the online approach adapts to non-stationary conditions as network load varies or LLM model updates change computational requirements.
+This capability validates WiLLM’s cross-layer APIs that provide real-time feedback enabling dynamic optimization.
+6. Related Works
+6.1. Wireless Systems for AI/ML Services
+Recent research has explored various approaches to supporting AI/ML workloads over wireless networks.
+Edge computing frameworks (Chen et al., 2024; Wu et al., 2020) position computational resources at network edges to reduce latency, but face fundamental limitations in supporting large language models due to power and thermal constraints.
+Cloud-edge collaborative architectures (Dong et al., 2024) attempt to balance computation between edge and cloud, yet struggle with the tight coupling between LLM inference and communication that our measurements reveal.
+Distributed inference systems (Xue et al., 2024; Mitra et al., 2024) partition LLM computation across multiple nodes.
+While these approaches address computational scalability, they introduce complex coordination challenges and remain largely theoretical without practical wireless implementations.
+In contrast, WiLLM provides a complete, operational system that researchers can immediately use for experimentation.
+Network slicing for AI services has received attention (Liu et al., 2024; Dandoush et al., 2024), but existing work focuses on simulation-based evaluation rather than real implementations.
+These studies identify the potential of slicing for service differentiation but lack the dynamic adaptation mechanisms necessary for LLM’s variable workloads.
+WiLLM’s Tree-Branch-Fruit architecture extends beyond static slicing to provide runtime adaptation based on observed performance.
+6.2. Open-Source Wireless Platforms
+Several open-source platforms provide wireless network implementations, each with distinct limitations for LLM service research.
+OpenAirInterface (Nikaein et al., 2014) offers comprehensive protocol implementations but requires deep wireless expertise and lacks LLM-specific features.
+srsRAN (Gomez-Miguelez et al., 2016) provides a lightweight alternative but similarly focuses on traditional communication services.
+Open5GS (Barbosa et al., 2024) implements core network functions without radio access capabilities, limiting end-to-end experimentation.
+Recent platforms like FlexRIC (Schmidt et al., 2021) and Janus (Foukas et al., 2023) introduce programmability for network intelligence but remain focused on traditional metrics and lack the cross-layer visibility essential for LLM services.
+Critically, all existing platforms require specialized UE hardware or protocol modifications for advanced features like network slicing, creating barriers for AI researchers.
+WiLLM addresses these limitations through our universal compatibility approach and comprehensive APIs that abstract wireless complexity.
+By building on OAI’s solid foundation while adding LLM-specific innovations, we provide a platform accessible to researchers without telecommunications expertise.
+6.3. LLM Deployment and Optimization
+The LLM deployment landscape includes various frameworks optimizing inference performance (NVIDIA, 2025; ggml-org, 2025), but these focus solely on computational efficiency without considering wireless communication constraints.
+Container-based solutions like Ollama (Ollama, 2025) simplify deployment but lack awareness of network conditions that significantly impact service quality in wireless environments.
+Recent work on LLM serving systems (Xu et al., 2025; Zhang et al., 2025) addresses challenges like long-context handling and request scheduling in datacenter environments.
+However, these solutions assume stable, high-bandwidth connectivity unavailable in wireless settings.
+WiLLM bridges this gap by providing network awared LLM serving that adapts to wireless channel variations and resource constraints.
+This survey of existing work underscores the absence of integrated solutions for wireless LLM service delivery.
+The dichotomy between wireless infrastructure platforms and LLM deployment frameworks has created a research void that WiLLM fills through its unified architecture, empirical foundation, and accessible implementation.
+Our contributions demonstrate that bridging these previously separate domains enables new research opportunities and practical applications impossible with existing fragmented approaches.
 7. Conclusion
-WiLLM is the first open-source wireless system for LLM services.
-Our “Tree-Branch-Fruit” slicing architecture, LLMs deployment in CNs, and cross-layer APIs address limitations in existing wireless infrastructure for LLM delivery.
-The implementation extends OpenAirInterface with components enabling dynamic resource allocation, universal device compatibility, and multi-slice scheduling while preserving backward compatibility.
-A dataset of 1.6+ million synchronized metrics provides empirical foundation for research, while our smart glasses case demonstrates viability in resource-constrained environments.
-WiLLM connects telecom evolution with AI, enabling wireless LLM service chain optimization.
-Future research efforts will focus on addressing the current system’s deficiencies by prioritizing improvements in system stability and fault recovery, ensuring strict compliance with protocols, and implementing security measures for cross-layer APIs.
+We presented WiLLM, the first open-source platform for wireless LLM service delivery.
+We identified three characteristics of LLM services over wireless network: bidirectional asymmetry, computation-communication coupling, and token unpredictability.
+WiLLM addresses these through the new slicing technology, universal compatibility, and cross-layer APIs, enabling AI researchers to experiment with real cellular networks without wireless expertise.
+The smart glasses case study validated our approach with stable 2-second multimodal responses.
+By democratizing wireless LLM research, WiLLM transforms a specialized domain into an accessible field for advancing mobile AI services.
 Appendix
 Appendix A Appendix Overview
-This appendix provides supplementary information, analysis, and technical details to support the main body of the paper.
-The contents are organized as follows: Appendix B presents a general comparison of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks.
+This appendix provides supplementary information, analysis, and technical details to support the main body of the paper. The contents are organized as follows:
+Appendix B presents a general comparison of WiLLM against existing open-source wireless communication systems and LLM deployment frameworks.
 This comparative analysis establishes WiLLM’s distinctive contributions within the telecommunications-AI convergence landscape.
@@
 More importantly, we conducted a causal analysis of the results shown in Figure.
-6 to Figure  10 of the dataset.
+3 to Figure  7 of the dataset.
 Appendix G introduces two novel benchmarking metrics - the LLM-Aware Resource Efficiency Index (LAREI) and the LLM Slice Efficiency Quotient (LSEQ).
@@
 By offering detailed comparative analysis, comprehensive system specifications, in-depth dataset insights, and novel performance evaluation frameworks, this supplementary material enables researchers to fully engage with the technical contributions of WiLLM and build upon its open-source foundation to advance the state of the art in wireless LLM service delivery.
-Appendix B Appendix for Section 2 (Related Work and Comparison): General Comparation of Systems
+Appendix B Appendix for Section 6 (Related Works): General Comparation of Systems
 Table 1. Comparison of Open Source Wireless Communication Systems and LLM deployment Systems
@@
 This comparative evaluation demonstrates how WiLLM transcends the limitations of existing systems through its specialized design for LLM service delivery in wireless environments.
-Appendix C Appendix for Section 2 (Related Work and Comparison): Detailed Technical Comparison of Systems
+Appendix C Appendix for Section 6 (Related Works): Detailed Technical Comparison of Systems
 Table 2. Technical Comparison of Wireless Systems
@@
 This empirical evidence confirms that WiLLM’s slicing architecture successfully enhances throughput efficiency, allowing more responsive performance scaling for LLM services while maintaining an appropriate baseline throughput level, which is crucial for consistent user experience in latency-sensitive LLM applications.
-F.1. Analysis of Uplink vs. Downlink Transmission Patterns (Figures 6, 7 and 8)
-The differential performance profiles across slice configurations illustrated in Figure 8 substantiate the efficacy of WiLLM’s “Tree-Branch-Fruit” slicing architecture, with differences primarily attributable to input modality variations.
-The pronounced disparity in inference time contribution (11.6% to 59.6%) stems fundamentally from the computational asymmetry between processing graphical inputs in uplink scenarios versus textual inputs in downlink scenarios.
-Image-based queries in uplink transmissions necessitate substantially higher computational overhead for visual feature extraction, semantic interpretation, and multimodal reasoning, consequently dominating the latency profile.
-Conversely, text-based inputs in downlink scenarios require comparatively less computational intensity, shifting the performance bottleneck toward transmission components (28.9% to 54.7% variation in transmission time).
-This input-dependent computational variance underscores the necessity of WiLLM’s specialized slicing approach, which adaptively allocates resources based on modality-specific requirements rather than uniform distribution models.
-The statistically significant performance differentials between slices confirm that WiLLM effectively maintains service isolation while accommodating the inherent computational asymmetry between visual and textual processing pipelines, validating the fundamental premise of modality-optimized “Fruit Slices” within the hierarchical resource management framework.
-It is important to note that our current implementation does not utilize specialized image generation models, which would likely introduce additional computational demands in the downlink scenario.
-Such models might shift the bottleneck toward inference processing rather than transmission, potentially altering the observed latency distribution patterns.
-Nevertheless, the empirical evidence clearly demonstrates that downlink image transmission presents a more significant bandwidth challenge than uplink image transmission, likely due to differences in compression efficiency and quality requirements between input and output images.
-F.2. Resource Allocation-Utilization Relationship Analysis (Figures 9 and 10)
-The non-linear correlation between PRB allocation patterns (Figure 9) and actual byte transmission volumes (Figure 10) reveals critical insights into resource utilization efficiency within wireless LLM service delivery.
-While WiLLM successfully implements differentiated PRB allocation conforming to slice policies, the observed discrepancy between allocated resources and realized throughput reflects complex cross-layer interdependencies spanning the entire system architecture.
-This phenomenon stems from multi-dimensional factors: (1) variable UE request patterns and buffer state fluctuations that create unpredictable traffic bursts; (2) server-side processing variability due to LLM inference time inconsistencies and computational resource contention; and (3) system execution mechanisms within the gNB layer, including scheduler implementation details, slice prioritization algorithms, and HARQ management policies.
-The visualization of slice-distinguished traffic with periodic transmission bursts demonstrates how these factors propagate across architectural boundaries, creating emergent system behaviors that cannot be predicted through isolated layer analysis.
-This empirical evidence substantiates the necessity of WiLLM’s cross-layer API framework, which enables coordinated adaptation across UE configuration, gNB resource allocation, and server processing prioritization based on observed end-to-end performance characteristics rather than theoretical layer-specific models—establishing a comprehensive feedback system essential for optimizing LLM service delivery in wireless environments.
-F.3. Dataset Collection Parameters
-F.3.1. UE Configuration Parameters
+F.1. Dataset Collection Parameters
+F.1.1. UE Configuration Parameters
 During data collection, we employed a configurable UE client with the following parameter ranges:
@@
 This approach simulated diverse real-world usage scenarios with varying resource constraints and quality requirements.
-F.3.2. Network Slicing Configuration
+F.1.2. Network Slicing Configuration
 The network slicing configuration used in our experiments consisted of three primary slices with different resource allocation parameters.
@@
 Figure.
-9, Figure 10 and Figure.
+6, Figure 7 and Figure.
 21 reveal critical insights into network slicing performance that challenge conventional resource allocation assumptions.